### v1

# Time-Correlated Video Bridge Matching

Abstract
Diffusion models excel in noise-to-data generation tasks, providing a mapping from a Gaussian distribution to a more complex data distribution. However they struggle to model translations between complex distributions, limiting their effectiveness in data-to-data tasks. While Bridge Matching (BM) models address this by finding the translation between data distributions, their application to time-correlated data sequences remains unexplored. This is a critical limitation for video generation and manipulation tasks, where maintaining temporal coherence is particularly important. To address this gap, we propose Time-Correlated Video Bridge Matching (TCVBM), a framework that extends BM to time-correlated data sequences in the video domain. TCVBM explicitly models inter-sequence dependencies within the diffusion bridge, directly incorporating temporal correlations into the sampling process. We compare our approach to classical methods based on bridge matching and diffusion models for three video-related tasks: frame interpolation, image-to-video generation, and video super-resolution. TCVBM achieves superior performance across multiple quantitative metrics, demonstrating enhanced generation quality and reconstruction fidelity.
1 INTRODUCTION
Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021b) have emerged as a powerful paradigm for generative modeling, achieving remarkable results in high-fidelity data synthesis (Saharia et al., 2022; Rombach et al., 2022; Arkhipkin et al., 2024; Labs, 2024). By iteratively denoising samples from a Gaussian distribution, these models excel at producing diverse and realistic outputs. However, despite their widespread adoption, diffusion models exhibit a critical limitation: they struggle to model translations between complex-structured data distributions. This shortcoming hinders their effectiveness in data-to-data tasks, where smooth bridging of the distributions is essential.
In contrast, Bridge Matching (BM) offers a principled solution to this problem by explicitly constructing a translation between arbitrary data distributions (Peluchetti, 2023b, a; Liu et al., 2022; Zhou et al., 2023). These methods learn vector field that connects source and target distributions, demonstrating strong performance in image-to-image tasks  (Shi et al., 2023; Liu et al., 2023). However, the translation between independent data sequences, the components of which are time-correlated, remains unaddressed. At the same time, this is exactly what video data is, where one data sample is a sequence of correlated frames. While existing methods assume to operate with video data samples without considering their internal structure  (Wang et al., 2025b), this omission can lead to a decrease in temporal consistency in the generated video.
Contribution. To address these issues, we propose Time-Correlated Video Bridge Matching (TCVBM), a novel framework that extends Bridge Matching to time-dependent video data. Unlike prior work, TCVBM explicitly takes into account inter-sequence dependencies, ensuring faithful translation between source and target distributions (Figure 1). We evaluate TCVBM on three video-related tasks: frame interpolation, image-to-video generation, and video super resolution. A comparison with classical approaches that do not take into account the correlation between frames, such as DDPM (Ho et al., 2020), DDIM (Song et al., 2021a), and Bridge Matching, demonstrates that TCVBM provides better temporal consistency and reconstruction quality.
2 RELATED WORKS
2.1 Bridge Models
Despite the widespread adoption of diffusion models in generative tasks (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021b), the Gaussian noise commonly used to initialize the generation process lacks meaningful structural information about the underlying data distribution. However, by matching the velocity fields with pre-defined transport maps, it is possible to construct powerful generative models that rival diffusion-based approaches (Lipman et al., 2023). Furthermore, Diffusion Bridge Models have been shown to provide a more flexible framework for generative tasks compared to standard diffusion processes (Bortoli et al., 2021; Liu et al., 2022). Particularly, bridge models have shown their superiority over diffusion baselines for image restoration (Delbracio and Milanfar, 2024), image translation (Zhou et al., 2024a), and image reconstruction (Liu et al., 2023). Despite this, the use of Bridge Matching to work with complex distributions of sequential data, the components of which are correlated, remains a practically unexplored topic. A recent attempt to extend Bridge Matching to image-to-video generation, overlooked the intrinsic temporal dependencies between video frames (Wang et al., 2025b). In our solution, we address this problem and design our own interpolant that explicitly accounts for the assumption of linear correlations among the components of the random vectors that represent video data samples.
2.2 Video Generation and Manipulation
Here, we provide a brief overview of the methods for some video-related tasks. Without loss of generality, we evaluate the applicability of our approach to the video domain using the following three tasks:
Frame Interpolation
aims to synthesize middle frames between two input images while ensuring temporal smoothness and content consistency. Traditional approaches often rely on optical flow estimation between keyframes (Niklaus and Liu, 2020; Lee et al., 2020; Park et al., 2021; Huang et al., 2022) or leverage convolutional features and attention maps for middle frame generation (Kalluri et al., 2023; Shi et al., 2022; Reda et al., 2022). Diffusion models were first used for video interpolation using bidirectional masking to synthesize middle frames (Voleti et al., 2022). Subsequent advances expanded this direction, including conditional diffusion-based generation (Danier et al., 2024), multi-stage cascaded refinement (Jain et al., 2024), adapted image-to-video diffusion models (Wang et al., 2025a), and techniques for large-motion generation (Shen et al., 2024). However, these methods do not explicitly model inter-frame correlations in video sequences. Similarly, event-based approaches (Chen et al., 2025; Zhang et al., 2025) incorporate additional motion cues but still rely on standard diffusion processes without accounting for temporal correlations.
Image-to-Video Generation
is a specific type of video generation, where an image is used as the input for a neural network, which then creates a video with consistent and semantically accurate movement. Diffusion models and recent advancements in generative architectures have taken this task to a new level, providing high-quality results (Zhang et al., 2023; Xing et al., 2023; Shi et al., 2024; Guo et al., 2023, 2024; Arkhipkin et al., 2025). However, the fundamental limitation of the diffusion process on the noise-to-data setup posed a risk of losing essential information that must be transferred from the image to the animated video. Different approaches tried to solve this without going beyond the diffusion process (Ren et al., 2024; Wu et al., 2023), until FrameBridge (Wang et al., 2025b) proposed an approach for image-to-video generation based on Bridge Matching. FrameBridge uses the input image as a prior instead of the Gaussian distribution used in the diffusion algorithm. Although this makes it reasonable to consider the image-to-video generation as a data-to-data task, this approach does not take into account the specifics of the internal video data structure, in fact adapting bridge matching algorithm from image to video domain. In this paper, we propose a method that explicitly assumes the correlation of frames with each other and uses this to obtain samples from bridge distribution.
Video Super Resolution
focuses on reconstructing high-resolution video frames from their degraded low-resolution versions. The application of diffusion models to this task is motivated by their strong generative prior, which is capable of synthesizing realistic details to overcome complex degradations. A central challenge is reconciling the stochastic nature of the diffusion process with the stringent requirement for temporal coherence across video frames. To resolve this, methods such as Upscale-A-Video (Zhou et al., 2024b) and MGLD (Yang et al., 2024) introduce explicit spatiotemporal constraints, using techniques like temporal layer integration and motion-guided loss optimization. At the same time, bridge-matching methods have been applied for super resolution in the image domain (Liu et al., 2023; Gushchin et al., 2025). In this work, we expand the application of these methods to video super resolution and propose a new approach that explicitly accounts for temporal coherence between video frames.
3 BACKGROUND ON BRIDGE MATCHING
We briefly review the Bridge Matching framework (Peluchetti, 2023b, a; Liu et al., 2022; Shi et al., 2023), which constructs diffusion processes for data translation, given a distribution of clean data $p(\mathbf{x}_{0})$ and corrupted data $p(\mathbf{x}_{T})$ on $\mathbb{R}^{D}$. The goal is to model a stochastic process that transitions from $\mathbf{x}_{0}\sim p(\mathbf{x}_{0})$ to $\mathbf{x}_{T}\sim p(\mathbf{x}_{T}\mid\mathbf{x}_{0})$, while incorporating a prior dynamics.
Consider a coupling $p(\mathbf{x}_{0},\mathbf{x}_{T})=p(\mathbf{x}_{0})\,p(\mathbf{x}_{T}\mid\mathbf{x}_{0})$, and let the prior process be defined by the stochastic differential equation (SDE):
$d\mathbf{x}_{t}=f(\mathbf{x}_{t},t)\,dt+g(t)\,d\mathbf{W}_{t},$ (1)
where $f(\mathbf{x}_{t},t)$ is a drift function, $g(t)$ is a time-dependent noise scale, and $\mathbf{W}_{t}$ is a standard Wiener process. For a fixed starting point $\mathbf{x}_{s}$, we denote the marginal of the prior process at time $t$ by $q(\mathbf{x}_{t}\mid\mathbf{x}_{s})$.
Bridge Distribution.
Given a pair $(\mathbf{x}_{0},\mathbf{x}_{t^{\prime}})$ from the prior, the posterior distribution of the process at time $t<t^{\prime}$, denoted as $q(\mathbf{x}_{t}\mid\mathbf{x}_{0},\mathbf{x}_{t^{\prime}}),$ is referred to as the bridge distribution. Using Bayes’ rule, it is expressed as:
$q(\mathbf{x}_{t}\mid\mathbf{x}_{0},\mathbf{x}_{t^{\prime}})=\frac{q(\mathbf{x}_{t^{\prime}}\mid\mathbf{x}_{t},\mathbf{x}_{0})\,q(\mathbf{x}_{t}\mid\mathbf{x}_{0})}{q(\mathbf{x}_{t^{\prime}}\mid\mathbf{x}_{0})}.$
Bridge Matching Dynamics.
Bridge Matching aims to construct a stochastic process that interpolates between $\mathbf{x}_{T}$ and $\mathbf{x}_{0}$ using a reverse-time SDE:
$d\mathbf{x}_{t}=\left\{f(\mathbf{x}_{t},t)-g^{2}(t)\,v^{*}(\mathbf{x}_{t},t)\right\}dt+g(t)\,d\bar{\mathbf{W}}_{t},$
where $\bar{\mathbf{W}}_{t}$ is a standard Wiener process under time reversal $t\leftarrow T-t$, and $dt$ denotes a negative infinitesimal timestep.
Learning Objective.
The drift function $v^{*}(\mathbf{x}_{t},t)$ is approximated using the following optimization objective:
$\min_{\phi}\mathbb{E}_{\mathbf{x}_{0},\mathbf{x}_{T},t}\left[\left\|v_{\phi}(\mathbf{x}_{t},t)-\nabla_{\mathbf{x}_{t}}\log q(\mathbf{x}_{t}\mid\mathbf{x}_{0})\right\|^{2}\right],$ (2)
where $\mathbf{x}_{0}\sim p(\mathbf{x}_{0})$, $\mathbf{x}_{T}\sim p(\mathbf{x}_{T}\mid\mathbf{x}_{0})$, and $\mathbf{x}_{t}\sim q(\mathbf{x}_{t}\mid\mathbf{x}_{0},\mathbf{x}_{T})$ is sampled from the bridge distribution. Time $t$ is sampled uniformly from the interval $[0,T]$.
This formulation provides a principled way to learn drift functions that guide the translation of corrupted data samples from $p(\mathbf{x}_{T})$ to clean data samples from $p(\mathbf{x}_{0})$ through learned diffusion processes.
4 METHOD
Figure 1: Comparison of the Bridge Matching and Time-Correlated Video Bridge Matching methods. The frames of the same video in the distributions $p(\mathbf{x}_{0})$ and $p(\mathbf{X}_{0})$ are indicated by dots of the same color. While Bridge Matching method makes the transition from one distribution to another without considering the relationship between frames, our approach treats one video as a single $\mathbf{X}_{0}$ data sequence and constructs the transition taking into account the internal correlation between video frames.
In this section, we introduce our proposed Time-Correlated Video Bridge Matching (TCVBM) method for modeling video data sequences. The core idea is to incorporate temporal correlations directly into the prior diffusion process, enabling better coherence and reconstruction of sequential data (Figure 1). We provide formal derivations and defer all proofs to the supplementary material section  A.
4.1 Time-Correlated Prior Process
We consider sequences of length $N$, represented as
${\mathbf{X}}=(\mathbf{x}^{1},\ldots,\mathbf{x}^{N}),$ (3)
where each $\mathbf{x}^{n}\in\mathbb{R}^{D}$ for $n=1,\ldots,N$. We aim to define a prior diffusion process that imposes an inductive bias toward temporal smoothness across elements.
Column-wise independence across features.
To model high-dimensional data efficiently, we assume that the $D$ feature dimensions evolve independently but share the same temporal dynamics. For each feature index $d=1,\ldots,D$, we define the time-dependent trajectory
$\mathbf{x}_{t}^{(d)}=\begin{bmatrix}x_{t}^{(d,1)}&\ldots&x_{t}^{(d,N)}\end{bmatrix}^{\top}\in\mathbb{R}^{N},$
which evolves by a stochastic differential equation (SDE):
$d\mathbf{x}_{t}^{(d)}=\left(\mathbf{A}\mathbf{x}_{t}^{(d)}+\mathbf{b}^{(d)}\right)dt+g(t)\,d\mathbf{W}_{t}^{(d)},$
where $\mathbf{A}\in\mathbb{R}^{N\times N}$ is a symmetric, invertible matrix encoding temporal correlations, $\mathbf{b}^{(d)}\in\mathbb{R}^{N}$ is a drift correction term, and $\mathbf{W}_{t}^{(d)}$ is a standard Wiener process.
Matrix form of the prior.
Equivalently, the full intermediate sequence $\mathbf{X}_{t}\in\mathbb{R}^{N\times D}$ evolves as:
$d\mathbf{X}_{t}=(\mathbf{A}\mathbf{X}_{t}+\mathbf{b})dt+g(t)\,d\mathbf{W}_{t},$
where $\mathbf{b}=[\mathbf{b}^{(1)}\,\ldots\,\mathbf{b}^{(D)}]\in\mathbb{R}^{N\times D}$, and $\mathbf{W}_{t}\in\mathbb{R}^{N\times D}$ is a matrix of independent Wiener processes across columns. In all expressions involving covariance and scores, formulas are applied column-wise. For example,
$\boldsymbol{\Sigma}_{t}^{-1}(\mathbf{X}_{t}-\boldsymbol{\mu}_{t})\in\mathbb{R}^{N\times D}$
denotes applying $\boldsymbol{\Sigma}_{t}^{-1}\in\mathbb{R}^{N\times N}$ independently to each column of $\mathbf{X}_{t}-\boldsymbol{\mu}_{t}$. Further, unless otherwise stated, we will assume a time-independent noise scale $g(t)=\sqrt{\epsilon}$ for simplicity.
We now derive the transition and bridge distributions for this prior, which are essential for bridge matching.
Proposition 1 (Correlated Process Score).
Let $\mathbf{X}_{t}$ follow the linear SDE:
$d\mathbf{X}_{t}=(\mathbf{A}\mathbf{X}_{t}+\mathbf{b})\,dt+\sqrt{\epsilon}\,d\mathbf{W}_{t},\quad\mathbf{X}_{0}\sim\delta_{\mathbf{X}_{0}},$ (4)
then the marginal distribution of $\mathbf{X}_{t}$ is Gaussian:
$q(\mathbf{X}_{t}|\mathbf{X}_{0})=\mathcal{N}(\mathbf{X}_{t}\mid\boldsymbol{\mu}_{t|0}(\mathbf{X}_{0}),\boldsymbol{\Sigma}_{t|0}),$ (5)
with
$\boldsymbol{\mu}_{t|0}(\mathbf{X}_{0})$ $=e^{\mathbf{A}t}\mathbf{X}_{0}+\left(e^{\mathbf{A}t}-I\right)\mathbf{A}^{-1}\mathbf{b},$ (6)
$\boldsymbol{\Sigma}_{t|0}$ $=\epsilon\,\frac{e^{2\mathbf{A}t}-I}{2}\mathbf{A}^{-1}.$ (7)
The score function is then given by
$\nabla_{\mathbf{X}_{t}}\log q(\mathbf{X}_{t}|\mathbf{X}_{0})=-\boldsymbol{\Sigma}_{t|0}^{-1}(\mathbf{X}_{t}-\boldsymbol{\mu}_{t|0}(\mathbf{X}_{0})).$ (8)
To perform bridge matching, one also needs to be able to sample from $q(\mathbf{X}_{t}|\mathbf{X}_{0},\mathbf{X}_{T})$.
Proposition 2 (Correlated Bridge Distribution).
Let $\mathbf{X}_{t}$ follow the same SDE as in Proposition 1. Then, given fixed endpoints $\mathbf{X}_{0}$ and $\mathbf{X}_{t^{\prime}}$, the posterior (bridge) distribution of $\mathbf{X}_{t}$ is Gaussian:
$q(\mathbf{X}_{t}|\mathbf{X}_{0},\mathbf{X}_{t^{\prime}})=\mathcal{N}(\mathbf{X}_{t}\mid\boldsymbol{\mu}_{t|0,t^{\prime}},\boldsymbol{\Sigma}_{t|0,t^{\prime}}),$ (9)
where
$\boldsymbol{\mu}_{t|0,t^{\prime}}$ $=\boldsymbol{\mu}_{t|0}(\mathbf{X}_{0})+\boldsymbol{\Sigma}_{t|0}\boldsymbol{\Sigma}_{t^{\prime}|0}^{-1}(\mathbf{X}_{t^{\prime}}-\boldsymbol{\mu}_{t^{\prime}|0}(\mathbf{X}_{0})),$ (10)
$\boldsymbol{\Sigma}_{t|0,t^{\prime}}$ $=\boldsymbol{\Sigma}_{t|0}-\boldsymbol{\Sigma}_{t|0}\boldsymbol{\Sigma}_{t^{\prime}|0}^{-1}\boldsymbol{\Sigma}_{t|0}.$ (11)
Together, Propositions 1 and 2 provide closed-form expressions required to implement bridge matching under the time-correlated prior.
4.2 Time-Correlated Video Bridge Matching
Training.
To train a bridge matching model, we follow the general framework of Bridge Matching described in section 3. We assume access to clean samples $\mathbf{X}_{0}\sim p_{0}(\mathbf{X}_{0})$, and a degradation process $p(\mathbf{X}_{T}|\mathbf{X}_{0})$, together forming a coupling $p(\mathbf{X}_{0},\mathbf{X}_{T})=p_{0}(\mathbf{X}_{0})p(\mathbf{X}_{T}|\mathbf{X}_{0})$.
We aim to minimize the squared error between the predicted score function $v_{\phi}(\mathbf{X}_{t},t)$ and the score of prior process $\nabla_{\mathbf{X}_{t}}\log p(\mathbf{X}_{t}|\mathbf{X}_{0})$, averaged over bridge samples $\mathbf{X}_{t}\sim p(\mathbf{X}_{t}|\mathbf{X}_{0},\mathbf{X}_{T})$:
$\min_{\phi}\mathbb{E}_{\mathbf{X}_{0},\mathbf{X}_{t},t}\left[\left\|v_{\phi}(\mathbf{X}_{t},t)+\boldsymbol{\Sigma}_{t|0}^{-1}(\mathbf{X}_{t}-\boldsymbol{\mu}_{t|0}(\mathbf{X}_{0}))\right\|^{2}\right],$ (12)
where $t\sim\text{Uniform}(0,T)$.
This objective can be simplified by reparameterizing the score function in terms of an intermediate predictor:
Proposition 3 (Reparameterization of the drift function).
The minimizer $v^{*}(\mathbf{X}_{t},t)$ of the objective (12) can be expressed as:
$v^{*}(\mathbf{X}_{t},t)=-\boldsymbol{\Sigma}_{t|0}^{-1}\left(\mathbf{X}_{t}-\boldsymbol{\mu}_{t|0}(\widehat{\mathbf{X}}_{0}^{*}(\mathbf{X}_{t},t))\right),$
where $\widehat{\mathbf{X}}_{0}^{*}(\mathbf{X}_{t},t)$ is the solution to the regression problem:
$\min_{\phi}\mathbb{E}_{\mathbf{X}_{0},\mathbf{X}_{t},t}\left[\|\widehat{\mathbf{X}}_{0}^{\phi}(\mathbf{X}_{t},t)-\mathbf{X}_{0}\|^{2}\right].$ (13)
Thus, learning the score function reduces to learning a predictor for the clean data $\mathbf{X}_{0}$.
We parameterize the predictor $\widehat{\mathbf{X}}_{0}^{\phi}(\mathbf{X}_{t},t)$ with a neural network and train it using the regression loss in (13). The training procedure is summarized in Algorithm 1.
Inference.
At inference time, given a corrupted sequence $\mathbf{X}_{T}\sim p(\mathbf{X}_{T})$, we perform iterative denoising using the learned predictor and the time-correlated bridge distribution. Given a schedule $0=t_{0}<t_{1}<\cdots<t_{N}=T$, we iteratively refine the estimate of $\mathbf{X}_{0}$ by sampling from posterior:
$\mathbf{X}_{t_{n-1}}\sim p(\mathbf{X}_{t_{n-1}}\mid\widehat{\mathbf{X}}_{0},\mathbf{X}_{t_{n}}),$
where $\widehat{\mathbf{X}}_{0}=\widehat{\mathbf{X}}_{0}^{\phi}(\mathbf{X}_{t_{n}},t_{n})$ obtained by using prediction of the trained model. This process is detailed in Algorithm 2.
Algorithm 1 Training
0: data from coupling $p_{0}(\mathbf{X}_{0})p_{T}(\mathbf{X}_{T}|\mathbf{X}_{0})$ and coefficients $\mathbf{A}$, $\mathbf{b}$ and $\epsilon$ for prior (1).
1:repeat
2:  $t\sim\mathcal{U}([0,1]),\mathbf{X}_{0}\sim p_{0}(\mathbf{X}_{0}),\mathbf{X}_{T}\sim p(\mathbf{X}_{T}\mid\mathbf{X}_{0})$
3:  $\mathbf{X}_{t}\sim q\bigl(\mathbf{X}_{t}\mid\mathbf{X}_{0},\mathbf{X}_{T}\bigr)$ (9)
4:  Take gradient descent step on $\mathbf{X}_{0}^{\phi}(\mathbf{X}_{t},t)$ (13)
5:until convergence
Algorithm 2 Inference
0: Input $\mathbf{X}_{T}\sim p_{T}(\mathbf{X}_{T})$, trained model $\widehat{\mathbf{X}}_{0}^{\phi}(\cdot,\cdot)$, time schedule $\{t_{n}\}_{n=0}^{N}$
1: Set $\mathbf{X}_{t_{N}}\leftarrow\mathbf{X}_{T}$
2:for $n=N$ to 11 do
3:  Predict $\widehat{\mathbf{X}}_{0}\leftarrow\widehat{\mathbf{X}}_{0}^{\phi}(\mathbf{X}_{t_{n}},t_{n})$
4:  Sample $\mathbf{X}_{t_{n-1}}\sim p(\mathbf{X}_{t_{n-1}}\mid\widehat{\mathbf{X}}_{0},\mathbf{X}_{t_{n}})$ using (9)
5:end for
6:return $\mathbf{X}_{0}=\mathbf{X}_{t_{0}}$
4.3 Choice of the Prior Process for Video Manipulation Tasks
To encourage smooth transitions between consecutive elements of the sequence, we define the prior matrix $\mathbf{A}$ of the size $N\times N$ with a tridiagonal structure:
$\mathbf{A}=\begin{bmatrix}-2&1&0&\cdots&0\\ 1&-2&1&\ddots&\vdots\\ 0&\ddots&\ddots&\ddots&0\\ \vdots&\ddots&1&-2&1\\ 0&\cdots&0&1&-2\end{bmatrix}$.
Here, $\mathbf{A}$ promotes temporal correlations across adjacent elements, and the per-element prior is:
$d\mathbf{x}_{t}^{n}=\big((\mathbf{x}_{t}^{n-1}-\mathbf{x}_{t}^{n})+(\mathbf{x}_{t}^{n+1}-\mathbf{x}_{t}^{n})\big)\,dt+\sqrt{\epsilon}\,d\mathbf{W}_{t},$ (14)
where $n=2,\dots,N-1$. This formulation naturally encourages a linear relationship between the elements of the sequence. From the point of view of this approximation, each frame $\mathbf{x}_{t}^{n}$ should remain close to the average of its neighbors $\mathbf{x}_{t}^{n-1}$ and $\mathbf{x}_{t}^{n+1}$. It is particularly well-suited for video-related tasks, where the frames of one video are correlated with each other, and the prior process enforces consistency and smoothness between them.
Depending on the video manipulation task, we suggest the following options for vector $\mathbf{b}$ and equation 14:
Frame Interpolation.
In this case, we consider the sequence of length $N+2$, represented as
${\widetilde{\mathbf{X}}}=(\mathbf{x}^{0},\mathbf{x}^{1},\ldots,\mathbf{x}^{N},\mathbf{x}^{N+1}),$
where the endpoints $\mathbf{x}^{0}$ and $\mathbf{x}^{N+1}$ are fixed as the initial and final frames of a video clip, between which it is necessary to make interpolation. The middle part of the video will be defined in the same way as in the equation 3:
${\mathbf{X}}=(\mathbf{x}^{1},\ldots,\mathbf{x}^{N}),$
and all statements for $\mathbf{X}_{t}$ from paragraphs 4.1 and 4.2 remain valid. Considering in equation 14 for all $t$ and $n=1,\ldots,N$ $\mathbf{x}_{t}^{0}=\mathbf{x}^{0}$ and $\mathbf{x}_{t}^{N+1}=\mathbf{x}^{N+1}$, we can define the vector $\mathbf{b}$ as:
$\mathbf{b}=\begin{bmatrix}\mathbf{x}^{0},0,\dots 0,\mathbf{x}^{N+1}\end{bmatrix}^{T},\quad\mathbf{b}\in\mathbb{R}^{N\times D},$
i.e. $\mathbf{b}$ enforces the boundary conditions from the known fixed endpoints $\mathbf{x}^{0}$ and $\mathbf{x}^{N+1}$.
Image-to-Video Generation.
This task corresponds to the case of one fixed left point of the video sequence, i.e.:
${\widetilde{\mathbf{X}}}=(\mathbf{x}^{0},\mathbf{x}^{1},\ldots,\mathbf{x}^{N}),\quad\widetilde{\mathbf{X}}\in\mathbb{R}^{(N+1)\times D}\\$
$\mathbf{b}=\begin{bmatrix}\mathbf{x}^{0},0,\dots 0\end{bmatrix}^{T},\quad\mathbf{b}\in\mathbb{R}^{N\times D},$
and in the equation 14 for all $n$ and $t$ $\mathbf{x}_{t}^{0}=\mathbf{x}^{0}$ and $\mathbf{x}_{t}^{N+1}=0$.
Video Super Resolution.
This is the simplest case, where $\widetilde{\mathbf{X}}=\mathbf{X}\in\mathbb{R}^{N\times D}$, i.e. the endpoints are not fixed, and the vector $\mathbf{b}$ is equal to zero.
Figure 2: Frame interpolation results. The first and last frames are fixed and come from the dataset. The task is to generate the intermediate 88 frames. TCVBM provides a smoother and more consistent transition from frame to frame and a better detalization.
Table 1: Frame interpolation quantitative results.
Metric FVD $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$
DDIM 33.609 0.105 15.800 0.760
DDPM 32.404 0.117 14.937 0.741
BM 34.315 0.079 17.103 0.794
TCVBM (ours) 30.542 0.077 17.280 0.813
Figure 3: Image-to-Video generation results. Based on the first frame, the methods need to generate the next 99 frames. As can be seen, TCVBM makes it possible to achieve better appearance of the generations and a lower quality degradation with increasing sequence length.
5 EXPERIMENTS
5.1 Implementation Details
To evaluate the performance of our new method and compare it to other approaches, we take our own small and simple U-Net model based on several residual blocks with $2D$ convolutions. The model has approximately $8.7$ million parameters. We conduct experiments on the MovingMNIST dataset (Srivastava et al., 2015) , which contains $10,000$ video sequences each $20$ frames long and showing $2$ moving digits in a $64\times 64$ resolution. For each experiment, we extract sub-sequences from the videos, consisting of $10$ consecutive frames, and concatenate them into a $10$-channel input for the neural network. We use $9,500$ training sequences and $500$ validation sequences, with a random split for each seed value. We train each model for $150,000$ iterations with a batch size of $128$ and an ema rate of $0.999$. We use the AdamW optimizer (Loshchilov and Hutter, 2019) with betas set to $0.9$ and $0.95$, a weight decay of $10^{-4}$, and a learning rate of $3\times 10^{-5}$. The number of steps in the reverse process for all methods is equal to $1000$. All experiments are conducted using a single NVIDIA Tesla A100100 GPU.
Table 2: Image-to-Video generation quantitative results.
Metric FVD $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$
DDIM 77.72 0.294 10.88 0.603
DDPM 75.86 0.311 10.72 0.595
BM 49.32 0.271 10.63 0.579
TCVBM (ours) 44.96 0.258 10.75 0.591
5.2 Experiments Roadmap
We compare TCVBM with three classical generative methods: DDPM (Ho et al., 2020), DDIM (Song et al., 2021a), and Bridge Matching with Brownian Bridge (BM) (Ibe, 2013, Chapter 9). For each method, we use the same setup, except for the loss function and sampling procedure in the forward and reverse processes, which are determined by the specific paradigm of each algorithm. Based on each of these four approaches, we train three models to solve the following tasks:
Frame Interpolation.
A sequence of $10$ frames is fed into the network as input, with the first and last frames from the dataset fixed. At the output, $8$ middle frames are generated, which are then compared with $8$ corresponding video frames from the dataset. The DDPM and DDIM methods aim to solve the generation problem for middle frames from random noise in the reverse process. For BM and TCVBM methods, we consider two ways to initialize the middle input frames: linear interpolation between the two fixed boundary frames, or noise samples from the Gaussian distribution $\mathcal{N}(\textbf{0},\mathbf{I})$. The second approach produced better results for BM and TCVBM, so the results presented here further are for noise-based initialization method. For more details, please see the Appendix C.1.
Image-to-Video Generation.
The task is to generate the remaining $9$ frames based on the first frame. We provide the first frame as input data, along with $9$ Gaussian samples from $\mathcal{N}(\textbf{0},\mathbf{I})$, or $9$ copies of the first frame. The second option demonstrated the best results for all methods, so further we consider this initialization variant. For more information, please see the Appendix C.2.
Video Super Resolution.
The network accepts $10$ low resolution frames, and $10$ frames with the dataset resolution are expected at the output. We consider starting with resolutions of $16\times 16$ or $32\times 32$ for all methods, but due to the simplicity of the task, we did not see any significant advantages for our method using the second option. Therefore, the further results are reported for the starting resolution of $16\times 16$. We also consider two initialization options: feeding the network with only low resolution data or low resolution data concatenated with Gaussian noise. Both approaches produced similar results, and here we use noise-free variant. More details can be found in the Appendix C.3.
Below, we present the results for the BM and TCVBM methods with $\epsilon$ equal to $0.1$ and a coefficient at $\mathbf{A}$ set to $1$ for the TCVBM model, i.e., $\widetilde{\mathbf{A}}=\alpha\mathbf{A}$ and $\widetilde{\mathbf{b}}=\alpha\mathbf{b}$, where $\alpha=1$. Additionally, we conduct a series of experiments to find optimal hyperparameters $\epsilon$ and $\alpha$. The results of these experiments can be found in Appendix D. Furthermore, we explored the dynamic correlation approach in which the coefficient $\alpha_{t}$ depends on time and increases as the reverse process progresses, i.e. when $t\rightarrow 0$. However, our experiments showed that this approach does not provide an advantage over TCVBM when the $\alpha$ is time-independent. The theoretical background and experimental findings are presented in Appendix E.
5.3 Results
Qualitative Comparison.
As can be seen in Figures 2, 3, and 4, TCVBM produces more consistent and accurate results compared to other generative methods. By considering the mutual correlation between frames, our approach ensures a smoother transition and information transfer between them, which enhances the overall method stability. Additional examples can be found in Appendix F.
Figure 4: Video super resolution results. The resolution is increased from $16\times 16$ to $64\times 64$. As can be seen, TCVBM outperforms other methods.
Table 3: Video super resolution quantitative results.
Metric FVD $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$
DDIM 334.700 0.514 17.307 0.613
DDPM 607.880 0.236 20.188 0.582
BM 53.516 0.020 22.642 0.954
TCVBM (ours) 59.491 0.020 22.670 0.970
Quantitative Evaluation.
We perform quantitative evaluation on $500$ videos from our MovingMNIST validation set using four metrics: FVD (Unterthiner et al., 2019), LPIPS (Zhang et al., 2018), PSNR, and SSIM. Tables 1, 2, and 3 present the metric values. As can be seen, in most cases TCVBM outperforms other approaches in terms of quantitative metrics. Additional results providing the standard deviation over launches with different random seeds are presented in the Appendix B.
6 DISCUSSION
Potential Impact.
The proposed Time-Correlated Video Bridge Matching (TCVBM) framework is designed for generative modeling and manipulation with sequential video data. Unlike traditional diffusion and bridge matching methods, which often ignore the intrinsic temporal structure of data, TCVBM explicitly models inter-sequence correlations. This principle are applicable not only to video but also to other types of sequences, such as audio signals or time series, where temporal consistency is important. The flexibility in defining prior process parameters, such as tridiagonal matrices for local correlations, allows for the method to be adapted to specific applications.
Limitations.
This work has several limitations. First, the reliance on a predefined tridiagonal matrix prior is likely insufficient for capturing the complex nature of temporal dependencies in real-world data, particularly long-range and non-linear dynamics. Consequently, the exploration of more sophisticated interpolants – potentially modeling correlations in a feature space via adapters rather than directly between frames-is a critical direction for future work. Second, the computational cost of the iterative sampling process precludes real-time application. Finally, while the proposed approach offers a novel theoretical framework, its empirical validation is currently limited to the MovingMNIST dataset. Broader evaluation on large-scale, real-world datasets, complemented by human evaluation, is essential to substantiate its practical effectiveness and robustness.
Supplementary Materials
Appendix A PROOF OF PROPOSITIONS
Proof of Proposition 1.
Consider the linear SDE
$d\mathbf{X}_{t}=(\mathbf{A}\mathbf{X}_{t}+\mathbf{b})dt+\sqrt{\epsilon}\,d\mathbf{W}_{t},\quad\mathbf{X}_{0}\sim\delta_{\mathbf{X}_{0}},$
with $\mathbf{A}\in\mathbb{R}^{D\times D}$ symmetric and invertible, $\mathbf{b}\in\mathbb{R}^{D}$, and a $D$-dimensional standard Wiener process $\mathbf{W}_{t}$.
Conditional mean. Let $\mathbf{\Phi}(t):=e^{\mathbf{A}t}$ and define $\mathbf{Y}_{t}:=(\mathbf{\Phi}(t))^{-1}\mathbf{X}_{t}=e^{-\mathbf{A}t}\mathbf{X}_{t}$ (note $\mathbf{Y}_{0}=\mathbf{X}_{0}$).
$d\mathbf{Y}_{t}=d\big(e^{-\mathbf{A}t}\mathbf{X}_{t}\big)=e^{-\mathbf{A}t}\,d\mathbf{X}_{t}+d(e^{-\mathbf{A}t})\,\mathbf{X}_{t}=$
$e^{-\mathbf{A}t}\big[(\mathbf{A}\mathbf{X}_{t}+\mathbf{b})\,dt+\sqrt{\epsilon}\,d\mathbf{W}_{t}\big]-\mathbf{A}e^{-\mathbf{A}t}\mathbf{X}_{t}\,dt=e^{-\mathbf{A}t}\mathbf{b}\,dt+\sqrt{\epsilon}\,e^{-\mathbf{A}t}d\mathbf{W}_{t},$
In the integral form:
$\mathbf{Y}_{t}=\mathbf{X}_{0}+\int_{0}^{t}e^{-\mathbf{A}s}\mathbf{b}\,ds+\sqrt{\epsilon}\int_{0}^{t}e^{-\mathbf{A}s}\,d\mathbf{W}_{s}.$
Multiplying by $\mathbf{\Phi}(t)=e^{\mathbf{A}t}$ yields:
$\mathbf{X}_{t}=e^{\mathbf{A}t}\mathbf{X}_{0}+\int_{0}^{t}e^{\mathbf{A}(t-s)}\mathbf{b}ds+\sqrt{\epsilon}\int_{0}^{t}e^{\mathbf{A}(t-s)}d\mathbf{W}_{s}.$ (15)
Hence, the conditional mean is:
$\boldsymbol{\mu}_{t|0}(\mathbf{X}_{0})=e^{\mathbf{A}t}\mathbf{X}_{0}+\Big(\!\int_{0}^{t}e^{\mathbf{A}(t-s)}ds\Big)\mathbf{b}=e^{\mathbf{A}t}X_{0}+\big(e^{\mathbf{A}t}-I\big)\mathbf{A}^{-1}\mathbf{b},$ (16)
Conditional variance.
$\mathbf{I}_{t}:=\int_{0}^{t}e^{A(t-s)}\,d\mathbf{W}_{s},\qquad\text{so that}\qquad\mathbf{X}_{t}-\boldsymbol{\mu}_{t|0}(\mathbf{X}_{0})=\sqrt{\epsilon}\,\mathbf{I}_{t}.$
$\mathrm{Cov}(\mathbf{X}_{t}\mid\mathbf{X}_{0})=\mathbb{E}\big[(\mathbf{X}_{t}-\boldsymbol{\mu}_{t|0})(\mathbf{X}_{t}-\boldsymbol{\mu}_{t|0})^{\top}\mid\mathbf{X}_{0}\big]=\epsilon\,\mathrm{Cov}(\mathbf{I}_{t}).$
In turn:
$\mathrm{Cov}(\mathbf{I}_{t})=\mathbb{E}\!\left[\left(\int_{0}^{t}e^{\mathbf{A}(t-s)}\,d\mathbf{W}_{s}\right)\left(\int_{0}^{t}e^{\mathbf{A}(t-r)}\,d\mathbf{W}_{r}\right)^{\!\top}\right].$
By Itô isometry:
$\mathbb{E}\!\left[\int_{0}^{t}G_{s}\,d\mathbf{W}_{s}\right]=0,\qquad\mathrm{Cov}\!\left(\int_{0}^{t}G_{s}\,d\mathbf{W}_{s},\int_{0}^{t}H_{s}\,dW_{s}\right)=\int_{0}^{t}G_{s}H_{s}^{\top}\,ds.$
Taking $G_{s}=H_{s}=e^{\mathbf{A}(t-s)}$ gives
$\mathrm{Cov}(\mathbf{I}_{t})=\int_{0}^{t}e^{\mathbf{A}(t-s)}\,e^{\mathbf{A}^{\top}(t-s)}\,ds.$
Because $\mathbf{A}$ is symmetric, $e^{\mathbf{A}^{\top}u}=e^{\mathbf{A}u}$, hence
$\mathrm{Cov}(\mathbf{I}_{t})=\int_{0}^{t}e^{2\mathbf{A}(t-s)}\,ds=\frac{1}{2}\big(e^{2\mathbf{A}t}-I\big)\mathbf{A}^{-1}.$
Consequently,
$\boldsymbol{\Sigma}_{t|0}=\mathrm{Cov}(\mathbf{X}_{t}\mid\mathbf{X}_{0})=\frac{\epsilon}{2}\big(e^{2\mathbf{A}t}-\mathbf{I}\big)\mathbf{A}^{-1}.$
Since $\mathbf{X}_{t}\mid\mathbf{X}_{0}$ is Gaussian with mean $\mu_{t|0}$ and covariance $\Sigma_{t|0}$, its score is
$\nabla_{\mathbf{X}_{t}}\log q(\mathbf{X}_{t}\mid\mathbf{X}_{0})=-\,\boldsymbol{\Sigma}_{t|0}^{-1}\big(\mathbf{X}_{t}-\boldsymbol{\mu}_{t|0}(\mathbf{X}_{0})\big).$
This completes the proof. ∎
Proof of Proposition 2.
Step 1: Joint law from the prior. From (15) and (16) in the proof of Proposition 1:
$\mathbf{X}_{u}=\boldsymbol{\mu}_{u|0}(\mathbf{X}_{0})+\sqrt{\epsilon}\int_{0}^{u}e^{\mathbf{A}(u-s)}d\mathbf{W}_{s},$
$\boldsymbol{\mu}_{u|0}(\mathbf{X}_{0})=e^{\mathbf{A}u}\mathbf{X}_{0}+\big(e^{\mathbf{A}u}-\mathbf{I}\big)\mathbf{A}^{-1}\mathbf{b}.$
Thus, conditionally on $\mathbf{X}_{0}$,
$\mathbb{E}[\mathbf{X}_{t}\mid\mathbf{X}_{0}]=\boldsymbol{\mu}_{t|0}(\mathbf{X}_{0}),\qquad\mathbb{E}[\mathbf{X}_{t^{\prime}}\mid\mathbf{X}_{0}]=\boldsymbol{\mu}_{t^{\prime}|0}(\mathbf{X}_{0}),$
$\boldsymbol{\Sigma}_{t|0}=\frac{\epsilon}{2}\big(e^{2\mathbf{A}t}-\mathbf{I}\big)\mathbf{A}^{-1},\qquad\boldsymbol{\Sigma}_{t^{\prime}|0}=\frac{\epsilon}{2}\big(e^{2\mathbf{A}t^{\prime}}-\mathbf{I}\big)\mathbf{A}^{-1}.$
For the cross-covariance, using Itô isometry and independence of increments, for $t<t^{\prime}$,
$\boldsymbol{\Sigma}_{t,t^{\prime}|0}=\mathrm{Cov}(\mathbf{X}_{t},\mathbf{X}_{t^{\prime}}\mid\mathbf{X}_{0})=\epsilon\int_{0}^{t}e^{\mathbf{A}(t-s)}\,e^{\mathbf{A}(t^{\prime}-s)}\,ds=\epsilon\int_{0}^{t}e^{\mathbf{A}(t+t^{\prime}-2s)}\,ds=\frac{\epsilon}{2}\,\mathbf{A}^{-1}\!\left(e^{\mathbf{A}(t+t^{\prime})}-e^{\mathbf{A}(t^{\prime}-t)}\right).$
Collecting blocks, we have the joint Gaussian (conditionally on $X_{0}$)
[…]
Step 2: Conditioning to obtain the bridge. For a joint Gaussian $\begin{bmatrix}x\\ y\end{bmatrix}$ with blocks $(\mu_{x},\mu_{y},\Sigma_{xx},\Sigma_{yy},\Sigma_{xy})$, the conditional $x\mid y$ is (Petersen et al., 2008, Section 8.1.3):
$x\mid y\sim\mathcal{N}\big(\mu_{x}+\Sigma_{xy}\Sigma_{yy}^{-1}(y-\mu_{y}),\Sigma_{xx}-\Sigma_{xy}\Sigma_{yy}^{-1}\Sigma_{yx}\big).$
Applying this with $x=\mathbf{X}_{t}$, $y=\mathbf{X}_{t^{\prime}}$ and the blocks above gives:
$\boldsymbol{\mu}_{t\mid 0,t^{\prime}}=\boldsymbol{\mu}_{t|0}(\mathbf{X}_{0})+\boldsymbol{\Sigma}_{t,t^{\prime}|0}\,\boldsymbol{\Sigma}_{t^{\prime}|0}^{-1}\!\big(\mathbf{X}_{t^{\prime}}-\boldsymbol{\mu}_{t^{\prime}|0}(\mathbf{X}_{0})\big),$
$\boldsymbol{\Sigma}_{t\mid 0,t^{\prime}}=\boldsymbol{\Sigma}_{t|0}-\boldsymbol{\Sigma}_{t,t^{\prime}|0}\,\boldsymbol{\Sigma}_{t^{\prime}|0}^{-1}\boldsymbol{\Sigma}_{t,t^{\prime}|0}.$
Here
$\boldsymbol{\Sigma}_{t|0}=\tfrac{\epsilon}{2}(e^{2\mathbf{A}t}-\mathbf{I})\mathbf{A}^{-1},\quad\boldsymbol{\Sigma}_{t^{\prime}|0}=\tfrac{\epsilon}{2}(e^{2\mathbf{A}t^{\prime}}-\mathbf{I})\mathbf{A}^{-1},\quad\boldsymbol{\Sigma}_{t,t^{\prime}|0}=\tfrac{\epsilon}{2}A^{-1}\big(e^{\mathbf{A}(t+t^{\prime})}-e^{\mathbf{A}(t^{\prime}-t)}\big).$
This completes the proof. ∎
Proof of Proposition 3.
Consider the following bijective reparameterization:
$v_{\phi}(\mathbf{X}_{t},t)=-\boldsymbol{\Sigma}_{t|0}^{-1}\left(\mathbf{X}_{t}-\boldsymbol{\mu}_{t|0}(\widehat{\mathbf{X}}_{0}^{\phi}(\mathbf{X}_{t},t))\right)$
and substitute it in the optimization problem:
$\min_{\phi}\mathbb{E}_{\mathbf{X}_{0},\mathbf{X}_{t},t}\left[\left\|v_{\phi}(\mathbf{X}_{t},t)+\boldsymbol{\Sigma}_{t|0}^{-1}(\mathbf{X}_{t}-\boldsymbol{\mu}_{t|0}(\mathbf{X}_{0}))\right\|^{2}\right],$
$\min_{\phi}\mathbb{E}_{\mathbf{X}_{0},\mathbf{X}_{t},t}\left[\left\|-\boldsymbol{\Sigma}_{t|0}^{-1}\left(\mathbf{X}_{t}-\boldsymbol{\mu}_{t|0}(\widehat{\mathbf{X}}_{0}^{\phi}(\mathbf{X}_{t},t))\right)+\boldsymbol{\Sigma}_{t|0}^{-1}(\mathbf{X}_{t}-\boldsymbol{\mu}_{t|0}(\mathbf{X}_{0}))\right\|^{2}\right]=$
$\min_{\phi}\mathbb{E}_{\mathbf{X}_{0},\mathbf{X}_{t},t}\left[\left\|\boldsymbol{\Sigma}_{t|0}^{-1}\left((\mathbf{X}_{t}-\boldsymbol{\mu}_{t|0}(\mathbf{X}_{0}))-\left(\mathbf{X}_{t}-\boldsymbol{\mu}_{t|0}(\widehat{\mathbf{X}}_{0}^{\phi}(\mathbf{X}_{t},t))\right)\right)\right\|^{2}\right]=$
$\min_{\phi}\mathbb{E}_{\mathbf{X}_{0},\mathbf{X}_{t},t}\left[\left\|\boldsymbol{\Sigma}_{t|0}^{-1}\left(\boldsymbol{\mu}_{t|0}(\widehat{\mathbf{X}}_{0}^{\phi}(\mathbf{X}_{t},t))-\boldsymbol{\mu}_{t|0}(\mathbf{X}_{0})\right)\right\|^{2}\right]=$
$\min_{\phi}\mathbb{E}_{\mathbf{X}_{0},\mathbf{X}_{t},t}\left[\left(\boldsymbol{\mu}_{t|0}(\widehat{\mathbf{X}}_{0}^{\phi}(\mathbf{X}_{t},t))-\boldsymbol{\mu}_{t|0}(\mathbf{X}_{0})\right)^{\top}(\boldsymbol{\Sigma}_{t|0}^{\top})^{-1}\boldsymbol{\Sigma}_{t|0}^{-1}\left(\boldsymbol{\mu}_{t|0}(\widehat{\mathbf{X}}_{0}^{\phi}(\mathbf{X}_{t},t))-\boldsymbol{\mu}_{t|0}(\mathbf{X}_{0})\right)\right]$
Taking the gradient of this objective with respect to $\phi$, we obtain:
$\mathbb{E}_{\mathbf{X}_{0},\mathbf{X}_{t},t}\left[2(\boldsymbol{\Sigma}_{t|0}^{\top})^{-1}\boldsymbol{\Sigma}_{t|0}^{-1}\left(\boldsymbol{\mu}_{t|0}(\widehat{\mathbf{X}}_{0}^{\phi}(\mathbf{X}_{t},t))-\boldsymbol{\mu}_{t|0}(\mathbf{X}_{0})\right)\right]=0$
Since $\boldsymbol{\Sigma}_{t|0}^{-1}$ is positive definite we can multiply by $\frac{1}{2}\boldsymbol{\Sigma}_{t|0}(\boldsymbol{\Sigma}_{t|0}^{\top})$ and get:
$\mathbb{E}_{\mathbf{X}_{0},\mathbf{X}_{t},t}\left[\boldsymbol{\mu}_{t|0}(\widehat{\mathbf{X}}_{0}^{\phi}(\mathbf{X}_{t},t))-\boldsymbol{\mu}_{t|0}(\mathbf{X}_{0})\right]=0$
Then for each $\mathbf{X}_{t}$ consider conditional mean:
$\mathbb{E}_{\mathbf{X}_{t},t}\left[\mathbb{E}_{\mathbf{X}_{0}|\mathbf{X}_{t}}\left[\boldsymbol{\mu}_{t|0}(\widehat{\mathbf{X}}_{0}^{\phi}(\mathbf{X}_{t},t))-\boldsymbol{\mu}_{t|0}(\mathbf{X}_{0})\right]\right]=0$
$\mathbb{E}_{\mathbf{X}_{0}|\mathbf{X}_{t}}\left[\boldsymbol{\mu}_{t|0}(\widehat{\mathbf{X}}_{0}^{\phi}(\mathbf{X}_{t},t))-\boldsymbol{\mu}_{t|0}(\mathbf{X}_{0})\right]=0$
From (16) we have:
$\mu_{t|0}(\mathbf{X}_{0})=e^{\mathbf{A}t}\mathbf{X}_{0}+\Big(\!\int_{0}^{t}e^{\mathbf{A}(t-s)}ds\Big)\mathbf{b}=e^{\mathbf{A}t}X_{0}+\big(e^{\mathbf{A}t}-I\big)\mathbf{A}^{-1}\mathbf{b},$
Then (note that $e^{\mathbf{A}t}$ is invertible and we can multiplu both sides on $e^{-At}$):
$\mathbb{E}_{\mathbf{X}_{0}|\mathbf{X}_{t}}\left[e^{\mathbf{A}t}\mathbf{X}_{0}^{\phi}(\mathbf{X}_{t},t)+\big(e^{\mathbf{A}t}-I\big)\mathbf{A}^{-1}\mathbf{b}-e^{\mathbf{A}t}\mathbf{X}_{0}+\big(e^{\mathbf{A}t}-I\big)\mathbf{A}^{-1}\mathbf{b}\right]=0$
$\mathbb{E}_{\mathbf{X}_{0}|\mathbf{X}_{t}}\left[e^{\mathbf{A}t}(\mathbf{X}_{0}^{\phi}(\mathbf{X}_{t},t)-\mathbf{X}_{0})\right]=0$
$\mathbb{E}_{\mathbf{X}_{0}|\mathbf{X}_{t}}\left[\mathbf{X}_{0}^{\phi}(\mathbf{X}_{t},t)-\mathbf{X}_{0}\right]=0$
$\mathbf{X}_{0}^{\phi}(\mathbf{X}_{t},t)=\mathbb{E}_{\mathbf{X}_{0}|\mathbf{X}_{t}}\left[\mathbf{X}_{0}\right]$
Hence, optimal $\mathbf{X}_{0}^{*}=\mathbb{E}_{\mathbf{X}_{0}|\mathbf{X}_{t}}\left[\mathbf{X}_{0}\right]$, which in turn is the minimizer of MSE problem:
$\min_{\phi}\mathbb{E}_{\mathbf{X}_{0},\mathbf{X}_{t},t}\left[\|\widehat{\mathbf{X}}_{0}^{\phi}(\mathbf{X}_{t},t)-\mathbf{X}_{0}\|^{2}\right].$
By substituting in to $v_{\phi}$ we have:
$v^{*}(\mathbf{X}_{t},t)=-\boldsymbol{\Sigma}_{t|0}^{-1}\left(\mathbf{X}_{t}-\boldsymbol{\mu}_{t|0}(\widehat{\mathbf{X}}_{0}^{*}(\mathbf{X}_{t},t))\right)$
This completes the proof. ∎
Appendix B ADDITIONAL QUANTITATIVE RESULTS
Here, we present the results of a quantitative comparison between DDPM, DDIM, Bridge Matching (BM), and our proposed method, TCVBM. The values of the standard deviation are provided, based on 3 runs of each method with different random seeds.
Table 4: Frame interpolation quantitative results with standard deviation. The best values in column are bold, second best values are underlined.
Metric FVD $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$
DDIM 34.664 $\pm$ 5.80 0.120 $\pm$ 0.070 15.843 $\pm$ 0.120 0.766 $\pm$ 0.011
DDPM 33.612 $\pm$ 1.494 0.107 $\pm$ 0.009 14.509 $\pm$ 0.427 0.714 $\pm$ 0.024
BM 34.766 $\pm$ 0.398 0.078 $\pm$ 0.001 17.265 $\pm$ 0.390 0.789 $\pm$ 0.005
TCVBM (ours) 31.491 $\pm$ 4.035 0.071 $\pm$ 0.019 17.451 $\pm$ 0.459 0.825 $\pm$ 0.044
Table 5: Image-to-Video generation quantitative results with standard deviation. The best values in column are bold, second best values are underlined.
Metric FVD $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$
DDIM 335.51 $\pm$ 241.12 0.402 $\pm$ 0.092 10.205 $\pm$ 0.514 0.513 $\pm$ 0.069
DDPM 250.52 $\pm$ 134.99 0.383 $\pm$ 0.054 10.333 $\pm$ 0.275 0.530 $\pm$ 0.046
BM 48.54 $\pm$ 0.56 0.268 $\pm$ 0.005 10.627 $\pm$ 0.053 0.582 $\pm$ 0.004
TCVBM (ours) 45.32 $\pm$ 0.91 0.260 $\pm$ 0.002 10.710 $\pm$ 0.028 0.589 $\pm$ 0.001
Table 6: Video super resolution quantitative results with standard deviation. The best values in column are bold, second best values are underlined.
Metric FVD $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$
DDIM 336.808 $\pm$ 5.175 0.520 $\pm$ 0.006 17.226 $\pm$ 0.103 0.600 $\pm$ 0.016
DDPM 614.288 $\pm$ 6.289 0.237 $\pm$ 0.001 20.152 $\pm$ 0.050 0.577 $\pm$ 0.004
BM 29.710 $\pm$ 20.683 0.026 $\pm$ 0.005 21.412 $\pm$ 1.040 0.941 $\pm$ 0.012
TCVBM (ours) 32.762 $\pm$ 23.153 0.029 $\pm$ 0.004 21.431 $\pm$ 1.419 0.941 $\pm$ 0.011
Appendix C INITIALIZATION EXPERIMENTS
In this section, we explore options for initializing or representing input data for bridge-based methods used in our work, namely for Bridge Matching with Browninan Bridge (BM) and Time-Correlated Video Bridge Matching (TCVBM). The interest in initialization is primarily due to the fact that presumably these methods are better suited to data-to-data translation tasks. The interest in exploring the effect of input data initialization on the quality of model performance stems primarily from the assumption that bridge-based approaches are better suited for data-to-data translation tasks.
C.1 Frame Interpolation
As input data for the network, we explored two options: filling in intermediate frames with Gaussian noise sampled from $\mathcal{N}(\textbf{0},\textbf{1})$ and using linear interpolation between fixed boundary frames $\mathbf{x}^{0}$ and $x^{N}$, i.e.:
$\mathbf{x}^{n}_{input}=\frac{n\mathbf{x}^{0}+(N-n)x^{N}}{N},\quad n=1,\dots,N-1.$
Table 7 compares the results of these initialization methods. As can be seen, filling intermediate frames with noise from a normal distribution produces better results than the initial linear interpolation.
Table 7: Analysis of the impact of initialization of input video data for bridge-based methods in the task of frame interpolation. Using noise from a normal distribution shows a clear advantage. The best values in column are bold, second best values are underlined.
Initialization method Method FVD $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$
Linear interpolation BM 34.804 0.109 15.439 0.756
TCVBM 31.944 0.092 16.275 0.782
Gaussian noise from $\mathcal{N}(\textbf{0},\textbf{1})$ BM 34.315 0.079 17.103 0.794
TCVBM 30.542 0.077 17.280 0.813
C.2 Image-to-Video Generation
Here we compare the following two types of initial initialization: duplicating the first frame in place of the frames to be generated (static video) and using random noise everywhere except the first frame. Static video initialization is superior to the noise option for both models (Table 8).
Table 8: Comparison of two types of initial initialization of input data for image-to-video generation.
Initialization method Method FVD $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$
Static video BM 49.32 0.271 10.63 0.579
TCVBM 44.96 0.258 10.75 0.591
Gaussian noise from $\mathcal{N}(\textbf{0},\textbf{1})$ BM 52.57 0.287 10.61 0.568
TCVBM 48.61 0.263 10.68 0.587
C.3 Video Super Resolution
Table 9: Comparison of two types of initial initialization of input data for video super resolution. We perform this comparison for low-resolution $32\times 32$.
Initialization method Method FVD $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$
Low-resolution video BM 9.501 0.014 24.888 0.972
TCVBM 9.496 0.012 24.970 0.973
Low-resolution video BM 9.556 0.012 24.892 0.973
concatenated with noise from $\mathcal{N}(\textbf{0},\textbf{1})$ TCVBM 9.646 0.012 24.988 0.973
Appendix D HYPERPARAMETERS SEARCHING
Here we compare different values of hyperparameters, namely the noise scaling value $\epsilon$ and the coefficient $\alpha$, which determines the degree of impact of the matrix $\mathbf{A}$ as $\widetilde{\mathbf{A}}:=\alpha\mathbf{A}$. Table 11 shows that in the case of frame interpolation, it is impossible to identify a clear dependence of the generation quality on the hyperparameters used, however, a sufficient amount of noise and not large values for the $\alpha$ coefficient are optimal. The results for video super resolution in Table 11 demonstrate that small values of $\epsilon$ and $\alpha$ are optimal for this task, which does not contradict the results for frame interpolation.
Table 10: The results of TCVBM training with various hyperparameters $\epsilon$ and $\alpha$ for frame interpolation. The best values in column are bold, second best values are underlined.
$\epsilon$ $\alpha$ FVD $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$
0.1 0.1 35.572 0.085 16.40 0.797
0.1 1 36.542 0.089 16.37 0.797
0.1 10 29.792 0.086 16.56 0.801
1 0.1 27.342 0.084 16.65 0.803
1 1 30.542 0.077 16.86 0.813
1 10 31.432 0.080 17.12 0.817
10 0.1 37.662 0.084 17.24 0.819
10 1 54.542 0.093 17.17 0.818
10 10 54.562 0.100 17.17 0.769
Table 11: The results of TCVBM training with various hyperparameters $\epsilon$ and $\alpha$ for video super resolution. We perform this comparison for low-resolution $32\times 32$.
$\epsilon$ $\alpha$ FVD $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$
0.1 0.1 9.496 0.012 24.970 0.973
0.1 1 10.413 0.013 24.358 0.969
1 0.1 13.226 0.019 23.004 0.959
1 1 15.023 0.022 22.120 0.949
10 0.1 18.458 0.033 20.085 0.921
10 1 20.746 0.039 19.283 0.906
Appendix E DYNAMICAL CORRELATION
E.1 Theory
Consider prior SDE with an additional multiplicative function $f(t)$ depending only on time $t$:
$d\mathbf{X}_{t}=f(t)(\mathbf{A}\mathbf{X}_{t}+\mathbf{b})dt+\sqrt{\epsilon}d\mathbf{W}_{t}.$
The derivation of formulas for this prior follows the same principles used for $f(t)=1$ in Appendix A.
Define:
$F(t)=\int_{0}^{t}f(\tau)\,d\tau,\qquad\mathbf{\Phi}(t)=e^{\mathbf{A}\,F(t)}.$
Then $\frac{d}{dt}\mathbf{\Phi}(t)=f(t)\,\mathbf{A}\,\mathbf{\Phi}(t)$, $\mathbf{\Phi}(0)=I$, and $\frac{d}{dt}\mathbf{\Phi}(t)^{-1}=-\,f(t)\,\mathbf{\Phi}(t)^{-1}\mathbf{A}$.
Conditional mean. Consider $\mathbf{Y}_{t}:=\mathbf{\Phi}(t)^{-1}\mathbf{X}_{t}$. By Itô’s rule:
$d\mathbf{Y}_{t}=\mathbf{\Phi}(t)^{-1}d\mathbf{X}_{t}+d(\mathbf{\Phi}(t)^{-1})\,\mathbf{X}_{t}=\mathbf{\Phi}(t)^{-1}f(t)(\mathbf{A}\mathbf{X}_{t}+\mathbf{b})\,dt+\sqrt{\epsilon}\,\mathbf{\Phi}(t)^{-1}d\mathbf{W}_{t}-f(t)\,\mathbf{\Phi}(t)^{-1}A\mathbf{X}_{t}\,dt=$
$=\mathbf{\Phi}(t)^{-1}f(t)\mathbf{b}\,dt+\sqrt{\epsilon}\,\mathbf{\Phi}(t)^{-1}d\mathbf{W}_{t}.$
Integrating from 0 to $t$ gives
$\mathbf{Y}_{t}=\mathbf{X}_{0}+\int_{0}^{t}\mathbf{\Phi}(s)^{-1}f(s)\,\mathbf{b}\,ds+\sqrt{\epsilon}\int_{0}^{t}\mathbf{\Phi}(s)^{-1}d\mathbf{W}_{s},$
and thus
$\mathbf{X}_{t}=\mathbf{\Phi}(t)\mathbf{X}_{0}+\mathbf{\Phi}(t)\!\int_{0}^{t}\mathbf{\Phi}(s)^{-1}f(s)\,\mathbf{b}\,ds+\sqrt{\epsilon}\,\mathbf{\Phi}(t)\!\int_{0}^{t}\mathbf{\Phi}(s)^{-1}d\mathbf{W}_{s}.$
Taking expectation and using $\mathbf{\Phi}(t)\mathbf{\Phi}(s)^{-1}=e^{\mathbf{A}(F(t)-F(s))}$,
$\boldsymbol{\mu}_{t|0}(\mathbf{X}_{0})=\mathbb{E}[\mathbf{X}_{t}|\mathbf{X}_{0}]=e^{\mathbf{A}F(t)}\mathbf{X}_{0}+\int_{0}^{t}e^{\mathbf{A}(F(t)-F(s))}f(s)\,\mathbf{b}\,ds.$
With the change of variables $u=F(s)$ (so du=f(s)dsdu=f(s)\,ds) this equals
$\int_{0}^{F(t)}e^{\mathbf{A}(F(t)-u)}\,du\,\mathbf{b}=\Big[-\,e^{\mathbf{A}(F(t)-u)}\mathbf{A}^{-1}\Big]_{u=0}^{F(t)}\mathbf{b}=\big(e^{\mathbf{A}F(t)}-I\big)\mathbf{A}^{-1}\mathbf{b}.$
Therefore
$\boldsymbol{\mu}_{t|0}(\mathbf{X}_{0})=e^{\mathbf{A}F(t)}\mathbf{X}_{0}+\big(e^{\mathbf{A}F(t)}-I\big)\mathbf{A}^{-1}\mathbf{b}.$
The mean $\mu_{t|0}(X_{0})$ of the process starting from a given $\mathbf{X}_{0}$ is given by:
$\boldsymbol{\mu}_{t|0}(\mathbf{X}_{0})=e^{\mathbf{A}F(t)}\mathbf{X}_{0}+\left(e^{\mathbf{A}F(t)}-I\right)\mathbf{A}^{-1}\mathbf{b}.$
Conditional variance. Let $\mathbf{Z}_{t}:=\mathbf{X}_{t}-\boldsymbol{\mu}_{t|0}$ be the centered process. Subtracting the mean SDE from the original SDE yields
$d\mathbf{Z}_{t}=f(t)\,\mathbf{A}\,\mathbf{Z}_{t}\,dt+\sqrt{\epsilon}\,d\mathbf{W}_{t}.$
Define the covariance $\boldsymbol{\Sigma}_{t|0}:=\mathbb{E}[\mathbf{Z}_{t}\mathbf{Z}_{t}^{\top}]$. Using Itô’s rule for $\mathbf{Z}_{t}\mathbf{Z}_{t}^{\top}$,
$d(\mathbf{Z}_{t}\mathbf{Z}_{t}^{\top})=(d\mathbf{Z}_{t})\mathbf{Z}_{t}^{\top}+\mathbf{Z}_{t}(d\mathbf{Z}_{t})^{\top}+d\mathbf{Z}_{t}(d\mathbf{Z}_{t})^{\top},$
where
$(d\mathbf{Z}_{t})\,\mathbf{Z}_{t}^{\top}=f(t)\mathbf{A}\mathbf{Z}_{t}\mathbf{Z}_{t}^{\top}\,dt+\sqrt{\epsilon}\,d\mathbf{W}_{t}\,\mathbf{Z}_{t}^{\top},$ (17)
$\mathbf{Z}_{t}\,(d\mathbf{Z}_{t})^{\top}=f(t)\mathbf{Z}_{t}\mathbf{Z}_{t}^{\top}\mathbf{A}^{\top}\,dt+\sqrt{\epsilon}\,\mathbf{Z}_{t}\,d\mathbf{W}_{t}^{\top},$ (18)
$(d\mathbf{Z}_{t})(d\mathbf{Z}_{t})^{\top}=\epsilon\,d\mathbf{W}_{t}d\mathbf{W}_{t}^{\top}=\epsilon\,I\,dt.$ (19)
Hence
$d(\mathbf{Z}_{t}\mathbf{Z}_{t}^{\top})=f(t)\big(\mathbf{A}\mathbf{Z}_{t}\mathbf{Z}_{t}^{\top}+\mathbf{Z}_{t}\mathbf{Z}_{t}^{\top}\mathbf{A}^{\top}\big)\,dt+\sqrt{\epsilon}\,d\mathbf{W}_{t}\,\mathbf{Z}_{t}^{\top}+\sqrt{\epsilon}\,\mathbf{Z}_{t}\,d\mathbf{W}_{t}^{\top}+\epsilon\,I\,dt.$
and taking expectations gives
$\frac{d}{dt}\boldsymbol{\Sigma}_{t|0}=f(t)\,\mathbf{A}\,\boldsymbol{\Sigma}_{t|0}+f(t)\,\boldsymbol{\Sigma}_{t|0}\mathbf{A}^{\top}+\epsilon\,I,\quad\boldsymbol{\Sigma}_{t|0}=0.$
To get the cross-covariance, we use:
$\mathbf{Z}_{t}=\sqrt{\epsilon}\int_{0}^{t}e^{\mathbf{A}(F(t)-F(s))}\,d\mathbf{W}_{s},\qquad\mathbf{Z}_{t^{\prime}}=\sqrt{\epsilon}\int_{0}^{t^{\prime}}e^{\mathbf{A}(F(t^{\prime})-F(u))}\,d\mathbf{W}_{u},$
The Itô isometry yields
$\boldsymbol{\Sigma}_{t,t^{\prime}|0}:=\mathrm{Cov}(\mathbf{Z}_{t},\mathbf{Z}_{t^{\prime}}^{\top})=\epsilon\int_{0}^{t}e^{\mathbf{A}(F(t)-F(s))}\,e^{\mathbf{A}(F(t^{\prime})-F(s))^{\top}}\,ds.$
For symmetric $\mathbf{A}$, $e^{(\cdot)^{\top}}=e^{(\cdot)}$ and, since these exponentials commute (all are $e^{\mathbf{A}(\cdot)}$),
$e^{\mathbf{A}(F(t)-F(s))}\,e^{\mathbf{A}(F(t^{\prime})-F(s))}=e^{\mathbf{A}(F(t^{\prime})-F(t))}\,e^{2\mathbf{A}(F(t)-F(s))}.$
Hence
$\boldsymbol{\Sigma}_{t,t^{\prime}|0}=e^{\mathbf{A}(F(t^{\prime})-F(t))}\,\epsilon\int_{0}^{t}e^{2\mathbf{A}(F(t)-F(s))}\,ds=e^{\mathbf{A}\,[F(t^{\prime})-F(t)]}\,\boldsymbol{\Sigma}_{t|0}.$
Summary. Thus, all three components: mean $\boldsymbol{\mu}_{t|0}(\mathbf{X}_{0})$, variance $\boldsymbol{\Sigma}_{t|0}$, and cross-covariance $\boldsymbol{\Sigma}_{t,t^{\prime}|0}$ are derived and can be used further in the same way as in the original case of $f(t)=1$.
E.2 Experimental Results
The continuous and time-decreasing function $f(t)$ sets the increasing values of the matrix $\mathbf{A}$ in the inverse diffusion process when $t\rightarrow 0$. Thus, the correlation of frames with each other in the generated video increases in the last steps of the inference. We conduct a series of experiments to investigate the effect of dynamic correlation and the choice of the function $f(t)$. We use the same experimental setting as described in Section 5.1 of the main paper, with a number of optimizer steps set to $120,000$. The frame interpolation results are presented in Table 12. As can be seen, our experiments do not demonstrate the advantages of using dynamic correlation compared to using constant values of the matrix $\mathbf{A}$. However, we observe significant differences in the quality of the results depending on the function $f(t)$. This demonstrates a complex structure of the relationship between video frames in the diffusion process, which our framework provides in the constant linear approximation.
Table 12: The results of the selection of the function $f(t)$, which determines the inverse dependence of the values of the matrix $\mathbf{A}$ on time $t$. The best values in column are bold, second best values are underlined.
$f(t)$ FVD $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$
1t1-t 49.569 0.115 13.938 0.752
12t1-2t 398.823 0.202 11.760 0.684
10.5t1-0.5t 427.942 0.226 12.063 0.620
$2\times(1-t)$ 409.495 0.195 12.239 0.624
$0.5\times(1-t)$ 269.899 0.182 12.297 0.676
$(1-t)^{2}$ 43.651 0.112 13.916 0.751
$e^{-t}$ 1910.002 0.973 3.600 0.002
$e^{-2t}$ 1216.250 0.629 7.243 0.119
$e^{-4t}$ 108.052 0.142 13.186 0.688
$e^{-8t}$ 50.078 0.128 13.321 0.726
Constant ($f(t)=\alpha=1$) 36.309 0.097 14.515 0.772
Appendix F ADDITIONAL VISUAL EXAMPLES
F.1 Frame Interpolation
Figure 5: Additional frame interpolation results.
F.2 Image-to-Video Generation
Figure 6: Additional image-to-video results.
F.3 Video Super Resolution
Figure 7: Additional video super resolution results. The resolution is increased from $16\times 16$ to $64\times 64$.