### v2

# Time-Correlated Video Bridge Matching

Abstract
Diffusion models excel in noise-to-data generation tasks, providing a mapping from a Gaussian distribution to a more complex data distribution. However they struggle to model translations between complex distributions, limiting their effectiveness in data-to-data tasks. While Bridge Matching models address this by finding the translation between data distributions, their application to time-correlated data sequences remains unexplored. This is a critical limitation for video generation and manipulation tasks, where maintaining temporal coherence is particularly important. To address this gap, we propose Time-Correlated Video Bridge Matching (TCVBM), a framework that extends BM to time-correlated data sequences in the video domain. TCVBM explicitly models inter-sequence dependencies within the diffusion bridge, directly incorporating temporal correlations into the sampling process. We compare our approach to classical methods based on bridge matching and diffusion models for three video-related tasks: frame interpolation, image-to-video generation, and video super-resolution. TCVBM achieves superior performance across multiple quantitative metrics, demonstrating enhanced generation quality and reconstruction fidelity.
1 Introduction
Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021b) have emerged as a powerful paradigm for generative modeling, achieving remarkable results in high-fidelity data synthesis (Saharia et al., 2022; Rombach et al., 2022; Arkhipkin et al., 2024; Labs, 2024; Arkhipkin et al., 2025a). By iteratively denoising samples from a Gaussian distribution, these models excel at producing diverse and realistic outputs. However, despite their widespread adoption, diffusion models exhibit a critical limitation: they struggle to model translations between complex-structured data distributions. This shortcoming hinders their effectiveness in data-to-data tasks, where smooth bridging of the distributions is essential.
In contrast, Bridge Matching (BM) offers a principled solution to this problem by explicitly constructing a translation between arbitrary data distributions (Peluchetti, 2023b; a; Liu et al., 2022; Zhou et al., 2023). These methods learn vector field that connects source and target distributions, demonstrating strong performance in image-to-image tasks  (Shi et al., 2023; Liu et al., 2023). However, the translation between independent data sequences, the components of which are time-correlated, remains unaddressed. At the same time, this is exactly what video data is, where one data sample is a sequence of correlated frames. While existing methods assume to operate with video data samples without considering their internal structure  (Wang et al., 2025b), this omission can lead to a decrease in temporal consistency in the generated video.
Contribution. To address these issues, we propose Time Correlated Video Bridge Matching (TCVBM), a novel framework that extends Bridge Matching to time-dependent video data. Unlike prior work, TCVBM explicitly takes into account inter-sequence dependencies, ensuring faithful translation between source and target distributions (Figure 1). We evaluate TCVBM on three video-related tasks: frame interpolation, image-to-video generation, and video super resolution. A comparison with classical approaches that do not take into account the correlation between frames, such as DDPM (Ho et al., 2020), DDIM (Song et al., 2021a), and Bridge Matching, demonstrates that TCVBM provides better temporal consistency and reconstruction quality.
2 Related Works
2.1 Bridge Models
Despite the success of diffusion models in generative tasks (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021b), their reliance on Gaussian noise as a prior lacks meaningful structural information about the data. In contrast, models that match velocity fields using pre-defined transport maps can achieve competitive performance (Lipman et al., 2023). Bridge Matching offers a particularly flexible framework, outperforming standard diffusion for tasks like image restoration, translation, and reconstruction (Delbracio and Milanfar, 2024; Zhou et al., 2024a; Liu et al., 2023). However, applying Bridge Matching to correlated sequential data, such as video, remains largely unexplored. A recent extension to image-to-video generation (Wang et al., 2025b) overlooked inherent temporal dependencies. Our approach addresses this by designing an interpolant that explicitly models the linear correlations between video frames.
2.2 Temporal Modeling for Video Data
Advances in video generation often adapt pre-trained image models by adding temporal modules like 3D convolutions or attention layers (Ho et al., 2022; Blattmann et al., 2023; Arkhipkin et al., 2023). This architectural specialization continues with Diffusion Transformer-based video models (Chen et al., 2023; Ma et al., 2025). A key challenge is the computational cost of attention, addressed by techniques such as sparse attention and adaptive masking (Zhang et al., 2025b; Xi et al., 2025; Mikhailov et al., 2025). However, modeling temporal dependencies within the generative dynamics and SDE prior structure remains unexplored. These prior works are orthogonal to our contribution and compatible with our approach.
2.3 Video Generation and Manipulation
Frame Interpolation
aims to synthesize middle frames from two inputs, ensuring smoothness and consistency. Traditional methods use optical flow (Niklaus and Liu, 2020; Lee et al., 2020; Park et al., 2021; Huang et al., 2022) or convolutional features with attention (Kalluri et al., 2023; Shi et al., 2022; Reda et al., 2022). Diffusion-based interpolation began with bidirectional masking (Voleti et al., 2022) and advanced through conditional generation (Danier et al., 2024), cascaded refinement (Jain et al., 2024), adapted image-to-video models (Wang et al., 2025a), and large-motion techniques (Shen et al., 2024). However, these methods lack explicit modeling of inter-frame correlations. Event-based approaches (Chen et al., 2025; Zhang et al., 2025a) add motion cues but also use standard diffusion without capturing temporal dependencies.
Image-to-Video Generation
creates a video from an input image, requiring consistent and accurate motion. Diffusion models have significantly advanced this field, producing high-quality results (Zhang et al., 2023; Xing et al., 2023; Shi et al., 2024; Guo et al., 2023; 2024; Arkhipkin et al., 2025b; a). However, the standard noise-to-data diffusion process risks losing essential information from the input image. While some approaches address this within the diffusion framework (Ren et al., 2024; Wu et al., 2023), FrameBridge (Wang et al., 2025b) reformulates the task using Bridge Matching, treating the input image as a prior instead of Gaussian noise. This frames generation as a data-to-data task but does not account for the inherent structure and correlations between video frames. Our proposed method explicitly models these inter-frame correlations to sample from the bridge distribution.
Video Super Resolution
reconstructs high-resolution videos from low-resolution inputs. Diffusion models are applied for their strong generative prior, which synthesizes realistic details to overcome degradation. A key challenge is ensuring temporal coherence within the inherently stochastic diffusion process. Recent methods address this by introducing explicit spatiotemporal constraints, such as temporal layer integration and motion-guided losses (Zhou et al., 2024b; Yang et al., 2024). Meanwhile, bridge-matching methods have shown promise for image super-resolution (Liu et al., 2023; Gushchin et al., 2025). In this work, we extend bridge matching to video super resolution, proposing a novel approach that explicitly models temporal coherence between frames.
3 Background on Bridge Matching
We briefly review the Bridge Matching framework (Peluchetti, 2023b; a; Liu et al., 2022; Shi et al., 2023), which constructs diffusion processes for data translation, given a distribution of clean data $p(\mathbf{x}_{0})$ and corrupted data $p(\mathbf{x}_{T})$ on $\mathbb{R}^{D}$. The goal is to model a stochastic process that transitions from $\mathbf{x}_{0}\sim p(\mathbf{x}_{0})$ to $\mathbf{x}_{T}\sim p(\mathbf{x}_{T}\mid\mathbf{x}_{0})$, while incorporating a prior dynamics.
Consider a coupling $p(\mathbf{x}_{0},\mathbf{x}_{T})=p(\mathbf{x}_{0})\,p(\mathbf{x}_{T}\mid\mathbf{x}_{0})$, and let the prior process be defined by the stochastic differential equation (SDE):
$d\mathbf{x}_{t}=f(\mathbf{x}_{t},t)\,dt+g(t)\,d\mathbf{W}_{t},$ (1)
where $f(\mathbf{x}_{t},t)$ is a drift function, $g(t)$ is a time-dependent noise scale, and $\mathbf{W}_{t}$ is a standard Wiener process. For a fixed starting point $\mathbf{x}_{s}$, we denote the marginal of the prior process at time $t$ by $q(\mathbf{x}_{t}\mid\mathbf{x}_{s})$.
Bridge Distribution.
Given a pair $(\mathbf{x}_{0},\mathbf{x}_{t^{\prime}})$ from the prior, the posterior distribution of the process at time $t<t^{\prime}$, denoted as $q(\mathbf{x}_{t}\mid\mathbf{x}_{0},\mathbf{x}_{t^{\prime}}),$ is referred to as the bridge distribution. Using Bayes’ rule, it is expressed as:
$q(\mathbf{x}_{t}\mid\mathbf{x}_{0},\mathbf{x}_{t^{\prime}})=\frac{q(\mathbf{x}_{t^{\prime}}\mid\mathbf{x}_{t},\mathbf{x}_{0})\,q(\mathbf{x}_{t}\mid\mathbf{x}_{0})}{q(\mathbf{x}_{t^{\prime}}\mid\mathbf{x}_{0})}.$
Bridge Matching Dynamics.
Bridge Matching aims to construct a stochastic process that interpolates between $\mathbf{x}_{T}$ and $\mathbf{x}_{0}$ using a reverse-time SDE:
$d\mathbf{x}_{t}=\left\{f(\mathbf{x}_{t},t)-g^{2}(t)\,v^{*}(\mathbf{x}_{t},t)\right\}dt+g(t)\,d\bar{\mathbf{W}}_{t},$
where $\bar{\mathbf{W}}_{t}$ is a standard Wiener process under time reversal $t\leftarrow T-t$, and $dt$ denotes a negative infinitesimal timestep.
Learning Objective.
The drift function $v^{*}(\mathbf{x}_{t},t)$ is approximated using the following optimization objective:
$\min_{\phi}\mathbb{E}_{\mathbf{x}_{0},\mathbf{x}_{T},t}\left[\left\|v_{\phi}(\mathbf{x}_{t},t)-\nabla_{\mathbf{x}_{t}}\log q(\mathbf{x}_{t}\mid\mathbf{x}_{0})\right\|^{2}\right],$ (2)
where $\mathbf{x}_{0}\sim p(\mathbf{x}_{0})$, $\mathbf{x}_{T}\sim p(\mathbf{x}_{T}\mid\mathbf{x}_{0})$, and $\mathbf{x}_{t}\sim q(\mathbf{x}_{t}\mid\mathbf{x}_{0},\mathbf{x}_{T})$ is sampled from the bridge distribution. Time $t$ is sampled uniformly from the interval $[0,T]$.
This formulation provides a principled way to learn drift functions that guide the translation of corrupted data samples from $p(\mathbf{x}_{T})$ to clean data samples from $p(\mathbf{x}_{0})$ through learned diffusion processes.
4 Method
Figure 1: Comparison of the Bridge Matching and Time-Correlated Video Bridge Matching methods. The frames of the same video in the distributions $p(\mathbf{x}_{0})$ and $p(\mathbf{X}_{0})$ are indicated by dots of the same color. While Bridge Matching method makes the transition from one distribution to another without considering the relationship between frames, our approach treats one video as a single $\mathbf{X}_{0}$ data sequence and constructs the transition taking into account the internal correlation between video frames.
In this section, we introduce our proposed Time-Correlated Video Bridge Matching (TCVBM) method for modeling video data sequences. The core idea is to incorporate temporal correlations directly into the prior diffusion process, enabling better coherence and reconstruction of sequential data (Figure 1). We provide formal derivations and defer all proofs to the supplementary material section  A.
4.1 Time-Correlated Prior Process
We consider sequences of length $N$, represented as
${\mathbf{X}}=(\mathbf{x}^{1},\ldots,\mathbf{x}^{N}),$ (3)
where each $\mathbf{x}^{n}\in\mathbb{R}^{D}$ for $n=1,\ldots,N$. We aim to define a prior diffusion process that imposes an inductive bias toward temporal smoothness across elements.
Column-wise independence across features.
To model high-dimensional data efficiently, we assume that the $D$ feature dimensions evolve independently but share the same temporal dynamics. For each feature index $d=1,\ldots,D$, we define the time-dependent trajectory
$\mathbf{x}_{t}^{(d)}=\begin{bmatrix}x_{t}^{(d,1)}&\ldots&x_{t}^{(d,N)}\end{bmatrix}^{\top}\in\mathbb{R}^{N},$
which evolves by a stochastic differential equation (SDE):
$d\mathbf{x}_{t}^{(d)}=\left(\mathbf{A}\mathbf{x}_{t}^{(d)}+\mathbf{b}^{(d)}\right)dt+g(t)\,d\mathbf{W}_{t}^{(d)},$
where $\mathbf{A}\in\mathbb{R}^{N\times N}$ is a symmetric, invertible matrix encoding temporal correlations, $\mathbf{b}^{(d)}\in\mathbb{R}^{N}$ is a drift correction term, and $\mathbf{W}_{t}^{(d)}$ is a standard Wiener process.
Matrix form of the prior.
Equivalently, the full intermediate sequence $\mathbf{X}_{t}\in\mathbb{R}^{N\times D}$ evolves as:
$d\mathbf{X}_{t}=(\mathbf{A}\mathbf{X}_{t}+\mathbf{b})dt+g(t)\,d\mathbf{W}_{t},$
where $\mathbf{b}=[\mathbf{b}^{(1)}\,\ldots\,\mathbf{b}^{(D)}]\in\mathbb{R}^{N\times D}$, and $\mathbf{W}_{t}\in\mathbb{R}^{N\times D}$ is a matrix of independent Wiener processes across columns. In all expressions involving covariance and scores, formulas are applied column-wise. For example,
$\bm{\Sigma}_{t}^{-1}(\mathbf{X}_{t}-\bm{\mu}_{t})\in\mathbb{R}^{N\times D}$
denotes applying $\bm{\Sigma}_{t}^{-1}\in\mathbb{R}^{N\times N}$ independently to each column of $\mathbf{X}_{t}-\bm{\mu}_{t}$. Further, unless otherwise stated, we will assume a time-independent noise scale $g(t)=\sqrt{\epsilon}$ for simplicity.
We now derive the transition and bridge distributions for this prior, which are essential for bridge matching.
Proposition 1 (Correlated Process Score).
Let $\mathbf{X}_{t}$ follow the linear SDE:
$d\mathbf{X}_{t}=(\mathbf{A}\mathbf{X}_{t}+\mathbf{b})\,dt+\sqrt{\epsilon}\,d\mathbf{W}_{t},\quad\mathbf{X}_{0}\sim\delta_{\mathbf{X}_{0}},$ (4)
then the marginal distribution of $\mathbf{X}_{t}$ is Gaussian:
$q(\mathbf{X}_{t}|\mathbf{X}_{0})=\mathcal{N}(\mathbf{X}_{t}\mid\bm{\mu}_{t|0}(\mathbf{X}_{0}),\bm{\Sigma}_{t|0}),$ (5)
with
$\bm{\mu}_{t|0}(\mathbf{X}_{0})$ $=e^{\mathbf{A}t}\mathbf{X}_{0}+\left(e^{\mathbf{A}t}-I\right)\mathbf{A}^{-1}\mathbf{b},$ (6)
$\bm{\Sigma}_{t|0}$ $=\epsilon\,\frac{e^{2\mathbf{A}t}-I}{2}\mathbf{A}^{-1}.$ (7)
The score function is then given by
$\nabla_{\mathbf{X}_{t}}\log q(\mathbf{X}_{t}|\mathbf{X}_{0})=-\bm{\Sigma}_{t|0}^{-1}(\mathbf{X}_{t}-\bm{\mu}_{t|0}(\mathbf{X}_{0})).$ (8)
To perform bridge matching, one also needs to be able to sample from $q(\mathbf{X}_{t}|\mathbf{X}_{0},\mathbf{X}_{T})$.
Proposition 2 (Correlated Bridge Distribution).
Let $\mathbf{X}_{t}$ follow the same SDE as in Proposition 1. Then, given fixed endpoints $\mathbf{X}_{0}$ and $\mathbf{X}_{t^{\prime}}$, the posterior (bridge) distribution of $\mathbf{X}_{t}$ is Gaussian:
$q(\mathbf{X}_{t}|\mathbf{X}_{0},\mathbf{X}_{t^{\prime}})=\mathcal{N}(\mathbf{X}_{t}\mid\bm{\mu}_{t|0,t^{\prime}},\bm{\Sigma}_{t|0,t^{\prime}}),$ (9)
where
$\bm{\mu}_{t|0,t^{\prime}}$ $=\bm{\mu}_{t|0}(\mathbf{X}_{0})+\bm{\Sigma}_{t|0}\bm{\Sigma}_{t^{\prime}|0}^{-1}(\mathbf{X}_{t^{\prime}}-\bm{\mu}_{t^{\prime}|0}(\mathbf{X}_{0})),$ (10)
$\bm{\Sigma}_{t|0,t^{\prime}}$ $=\bm{\Sigma}_{t|0}-\bm{\Sigma}_{t|0}\bm{\Sigma}_{t^{\prime}|0}^{-1}\bm{\Sigma}_{t|0}.$ (11)
Together, Propositions 1 and 2 provide closed-form expressions required to implement bridge matching under the time-correlated prior.
4.2 Time-Correlated Video Bridge Matching
Training.
To train a bridge matching model, we follow the general framework of Bridge Matching described in section 3. We assume access to clean samples $\mathbf{X}_{0}\sim p_{0}(\mathbf{X}_{0})$, and a degradation process $p(\mathbf{X}_{T}|\mathbf{X}_{0})$, together forming a coupling $p(\mathbf{X}_{0},\mathbf{X}_{T})=p_{0}(\mathbf{X}_{0})p(\mathbf{X}_{T}|\mathbf{X}_{0})$.
We aim to minimize the squared error between the predicted score function $v_{\phi}(\mathbf{X}_{t},t)$ and the score of prior process $\nabla_{\mathbf{X}_{t}}\log p(\mathbf{X}_{t}|\mathbf{X}_{0})$, averaged over bridge samples $\mathbf{X}_{t}\sim p(\mathbf{X}_{t}|\mathbf{X}_{0},\mathbf{X}_{T})$:
$\min_{\phi}\mathbb{E}_{\mathbf{X}_{0},\mathbf{X}_{t},t}\left[\left\|v_{\phi}(\mathbf{X}_{t},t)+\bm{\Sigma}_{t|0}^{-1}(\mathbf{X}_{t}-\bm{\mu}_{t|0}(\mathbf{X}_{0}))\right\|^{2}\right],$ (12)
where $t\sim\text{Uniform}(0,T)$.
This objective can be simplified by reparameterizing the score function in terms of an intermediate predictor:
Proposition 3 (Reparameterization of the drift function).
The minimizer $v^{*}(\mathbf{X}_{t},t)$ of the objective equation 12 can be expressed as:
$v^{*}(\mathbf{X}_{t},t)=-\bm{\Sigma}_{t|0}^{-1}\left(\mathbf{X}_{t}-\bm{\mu}_{t|0}(\widehat{\mathbf{X}}_{0}^{*}(\mathbf{X}_{t},t))\right),$
where $\widehat{\mathbf{X}}_{0}^{*}(\mathbf{X}_{t},t)$ is the solution to the regression problem:
$\min_{\phi}\mathbb{E}_{\mathbf{X}_{0},\mathbf{X}_{t},t}\left[\|\widehat{\mathbf{X}}_{0}^{\phi}(\mathbf{X}_{t},t)-\mathbf{X}_{0}\|^{2}\right].$ (13)
Thus, learning the score function reduces to learning a predictor for the clean data $\mathbf{X}_{0}$.
We parameterize the predictor $\widehat{\mathbf{X}}_{0}^{\phi}(\mathbf{X}_{t},t)$ with a neural network and train it using the regression loss in equation 13. The training procedure is summarized in Algorithm 1.
Inference.
At inference time, given a corrupted sequence $\mathbf{X}_{T}\sim p(\mathbf{X}_{T})$, we perform iterative denoising using the learned predictor and the time-correlated bridge distribution. Given a schedule $0=t_{0}<t_{1}<\cdots<t_{N}=T$, we iteratively refine the estimate of $\mathbf{X}_{0}$ by sampling from posterior:
$\mathbf{X}_{t_{n-1}}\sim p(\mathbf{X}_{t_{n-1}}\mid\widehat{\mathbf{X}}_{0},\mathbf{X}_{t_{n}}),$
where $\widehat{\mathbf{X}}_{0}=\widehat{\mathbf{X}}_{0}^{\phi}(\mathbf{X}_{t_{n}},t_{n})$ obtained by using prediction of the trained model. This process is detailed in Algorithm 2.
Algorithm 1 Training
0: data from coupling $p_{0}(\mathbf{X}_{0})p_{T}(\mathbf{X}_{T}|\mathbf{X}_{0})$ and coefficients $\mathbf{A}$, $\mathbf{b}$ and $\epsilon$ for prior equation 1.
1:repeat
2:  $t\sim\mathcal{U}([0,1]),\mathbf{X}_{0}\sim p_{0}(\mathbf{X}_{0}),\mathbf{X}_{T}\sim p(\mathbf{X}_{T}\mid\mathbf{X}_{0})$
3:  $\mathbf{X}_{t}\sim q\bigl(\mathbf{X}_{t}\mid\mathbf{X}_{0},\mathbf{X}_{T}\bigr)$ equation 9
4:  Take gradient descent step on $\mathbf{X}_{0}^{\phi}(\mathbf{X}_{t},t)$ equation 13
5:until convergence
Algorithm 2 Inference
0: Input $\mathbf{X}_{T}\sim p_{T}(\mathbf{X}_{T})$, trained model $\widehat{\mathbf{X}}_{0}^{\phi}(\cdot,\cdot)$, time schedule $\{t_{n}\}_{n=0}^{N}$
1: Set $\mathbf{X}_{t_{N}}\leftarrow\mathbf{X}_{T}$
2:for $n=N$ to 11 do
3:  Predict $\widehat{\mathbf{X}}_{0}\leftarrow\widehat{\mathbf{X}}_{0}^{\phi}(\mathbf{X}_{t_{n}},t_{n})$
4:  Sample $\mathbf{X}_{t_{n-1}}\sim p(\mathbf{X}_{t_{n-1}}\mid\widehat{\mathbf{X}}_{0},\mathbf{X}_{t_{n}})$ using equation 9
5:end for
6:return $\mathbf{X}_{0}=\mathbf{X}_{t_{0}}$
4.3 Choice of the Prior Process for Video Manipulation Tasks
To encourage smooth transitions between consecutive elements of the sequence, we define the prior matrix $\mathbf{A}$ of the size $N\times N$ with a tridiagonal structure:
$\mathbf{A}=\begin{bmatrix}-2&1&0&\cdots&0\\ 1&-2&1&\ddots&\vdots\\ 0&\ddots&\ddots&\ddots&0\\ \vdots&\ddots&1&-2&1\\ 0&\cdots&0&1&-2\end{bmatrix}$.
Here, $\mathbf{A}$ promotes temporal correlations across adjacent elements, and the per-element prior is:
$d\mathbf{x}_{t}^{n}=\big((\mathbf{x}_{t}^{n-1}-\mathbf{x}_{t}^{n})+(\mathbf{x}_{t}^{n+1}-\mathbf{x}_{t}^{n})\big)\,dt+\sqrt{\epsilon}\,d\mathbf{W}_{t},$ (14)
where $n=2,\dots,N-1$. This formulation naturally encourages a linear relationship between the elements of the sequence. From the point of view of this approximation, each frame $\mathbf{x}_{t}^{n}$ should remain close to the average of its neighbors $\mathbf{x}_{t}^{n-1}$ and $\mathbf{x}_{t}^{n+1}$. It is particularly well-suited for video-related tasks, where the frames of one video are correlated with each other, and the prior process enforces consistency and smoothness between them.
Depending on the video manipulation task, we suggest the following options for vector $\mathbf{b}$ and equation 14:
Frame Interpolation.
In this case, we consider the sequence of length $N+2$, represented as
${\widetilde{\mathbf{X}}}=(\mathbf{x}^{0},\mathbf{x}^{1},\ldots,\mathbf{x}^{N},\mathbf{x}^{N+1}),$
where the endpoints $\mathbf{x}^{0}$ and $\mathbf{x}^{N+1}$ are fixed as the initial and final frames of a video clip, between which it is necessary to make interpolation. The middle part of the video will be defined in the same way as in the equation 3:
${\mathbf{X}}=(\mathbf{x}^{1},\ldots,\mathbf{x}^{N}),$
and all statements for $\mathbf{X}_{t}$ from paragraphs 4.1 and 4.2 remain valid. Considering in equation 14 for all $t$ and $n=1,\ldots,N$ $\mathbf{x}_{t}^{0}=\mathbf{x}^{0}$ and $\mathbf{x}_{t}^{N+1}=\mathbf{x}^{N+1}$, we can define the vector $\mathbf{b}$ as:
$\mathbf{b}=\begin{bmatrix}\mathbf{x}^{0},0,\dots 0,\mathbf{x}^{N+1}\end{bmatrix}^{T},\quad\mathbf{b}\in\mathbb{R}^{N\times D},$
i.e. $\mathbf{b}$ enforces the boundary conditions from the known fixed endpoints $\mathbf{x}^{0}$ and $\mathbf{x}^{N+1}$.
Image-to-Video Generation.
This task corresponds to the case of one fixed left point of the video sequence, i.e.:
${\widetilde{\mathbf{X}}}=(\mathbf{x}^{0},\mathbf{x}^{1},\ldots,\mathbf{x}^{N}),\quad\widetilde{\mathbf{X}}\in\mathbb{R}^{(N+1)\times D}\\$
$\mathbf{b}=\begin{bmatrix}\mathbf{x}^{0},0,\dots 0\end{bmatrix}^{T},\quad\mathbf{b}\in\mathbb{R}^{N\times D},$
and in the equation 14 for all $n$ and $t$ $\mathbf{x}_{t}^{0}=\mathbf{x}^{0}$ and $\mathbf{x}_{t}^{N+1}=0$.
Video Super Resolution.
This is the simplest case, where $\widetilde{\mathbf{X}}=\mathbf{X}\in\mathbb{R}^{N\times D}$, i.e. the endpoints are not fixed, and the vector $\mathbf{b}$ is equal to zero.
5 Experiments
5.1 Experiments Roadmap
We compare TCVBM with DDPM (Ho et al., 2020), DDIM (Song et al., 2021a), and Bridge Matching (BM) (Ibe, 2013, Chapter 9). Experiments use the MovingMNIST dataset (Srivastava et al., 2015), which contains $10,000$ video sequences each $20$ frames long and showing $2$ moving digits in a $64\times 64$ resolution. Implementation details can be found in the Appendix B. Using each approach, we train three models for the following tasks:
Frame Interpolation.
The network inputs a sequence of $10$ frames, with the first and last frames fixed from the dataset. It generates $8$ middle frames, which are compared to the ground truth. DDPM and DDIM generate these frames from random noise. For BM and TCVBM, we tested initializing the middle frames via linear interpolation or Gaussian noise from $\mathcal{N}(\textbf{0},\mathbf{I})$; noise yielded better results and is used further. Details can be found in the Appendix D.1.
Image-to-Video Generation.
The task is to generate the remaining $9$ frames based on the first frame. We provide the first frame as input data, along with $9$ Gaussian samples from $\mathcal{N}(\textbf{0},\mathbf{I})$, or $9$ copies of the first frame. The second option demonstrated the best results for all methods, so further we consider this initialization variant. For more information, please see the Appendix D.2.
(a) Frame interpolation. The first and last frames are fixed and come from the dataset. The task is to generate the intermediate 88 frames. TCVBM provides a smoother and more consistent transition from frame to frame and a better detalization.
(b) Image-to-Video generation. Based on the first frame, the method generate the next 99 frames. As can be seen, TCVBM makes it possible to achieve better appearance of the generations and a lower quality degradation with increasing sequence length.
(c) Video super resolution. The resolution is increased from $16\times 16$ to $64\times 64$. As can be seen, TCVBM outperforms other methods.
Figure 2: The results on the MovingMNIST dataset Srivastava et al. (2015).
Video Super Resolution.
The network inputs $10$ low-resolution frames and outputs $10$ frames at the dataset resolution. We evaluated starting resolutions of $16\times 16$ and $32\times 32$; as the simpler task showed no significant advantage for our method at $32\times 32$, we report results for $16\times 16$. We also tested initialization with low-resolution data alone versus data concatenated with Gaussian noise; both performed similarly, and we proceed with the noise-free version. See details in Appendix D.3.
We present results for BM and TCVBM with $\epsilon=0.1$ and $\alpha=1$ for TCVBM (i.e., $\widetilde{\mathbf{A}}=\mathbf{A}$, $\widetilde{\mathbf{b}}=\mathbf{b}$) for all three tasks. Experiments optimizing hyperparameters $\epsilon$ and $\alpha$ are in Appendix E. Computational complexity analysis is in Appendix F. We also explored a dynamic correlation approach where $\alpha_{t}$ increases as $t\to 0$, but found no advantage over constant $\alpha$ (see Appendix G).
5.2 Results
Qualitative Comparison.
As shown in Figure 2, TCVBM yields more consistent and accurate results than other methods. By modeling mutual correlation between frames, it ensures smoother transitions and better information transfer, improving overall stability. More examples can be found in the Appendix H.
Table 1: Frame interpolation quantitative results.
Metric FVD $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$
DDIM 33.609 0.105 15.800 0.760
DDPM 32.404 0.117 14.937 0.741
BM 34.315 0.079 17.103 0.794
TCVBM (ours) 30.542 0.077 17.280 0.813
Table 2: Image-to-Video generation quantitative results.
Metric FVD $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$
DDIM 77.72 0.294 10.88 0.603
DDPM 75.86 0.311 10.72 0.595
BM 49.32 0.271 10.63 0.579
TCVBM (ours) 44.96 0.258 10.75 0.591
Table 3: Video super resolution quantitative results.
Metric FVD $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$
DDIM 334.700 0.514 17.307 0.613
DDPM 607.880 0.236 20.188 0.582
BM 53.516 0.020 22.642 0.954
TCVBM (ours) 59.491 0.020 22.670 0.970
Quantitative Evaluation.
We perform quantitative evaluation on $500$ videos from our MovingMNIST validation set using FVD (Unterthiner et al., 2019), LPIPS (Zhang et al., 2018), PSNR, and SSIM. As shown in Tables 1, 2, and 3, TCVBM outperforms other methods on most metrics. Additional results, including standard deviations across random seeds, are in Appendix C.
6 Discussion
Potential Impact.
The proposed Time-Correlated Video Bridge Matching (TCVBM) framework is designed for generative modeling and manipulation with sequential video data. Unlike traditional diffusion and bridge matching methods, which often ignore the intrinsic temporal structure of data, TCVBM explicitly models inter-sequence correlations. This principle are applicable not only to video but also to other types of sequences, such as audio signals or time series, where temporal consistency is important. The flexibility in defining prior process parameters, such as tridiagonal matrices for local correlations, allows for the method to be adapted to specific applications.
Limitations.
This work has several limitations. First, the predefined tridiagonal prior may be insufficient for complex, long-range, or non-linear temporal dependencies in real-world data; future work should explore more sophisticated interpolants, such as those modeling correlations in feature space. Second, due to the computational cost, empirical validation is currently limited to MovingMNIST dataset. Broader evaluation on large-scale real-world datasets and human assessment are needed to confirm its practical effectiveness and robustness.
Appendix A Proof of Propositions
Proof of Proposition 1.
Consider the linear SDE
$d\mathbf{X}_{t}=(\mathbf{A}\mathbf{X}_{t}+\mathbf{b})dt+\sqrt{\epsilon}\,d\mathbf{W}_{t},\quad\mathbf{X}_{0}\sim\delta_{\mathbf{X}_{0}},$
with $\mathbf{A}\in\mathbb{R}^{D\times D}$ symmetric and invertible, $\mathbf{b}\in\mathbb{R}^{D}$, and a $D$-dimensional standard Wiener process $\mathbf{W}_{t}$.
Conditional mean. Let $\mathbf{\Phi}(t):=e^{\mathbf{A}t}$ and define $\mathbf{Y}_{t}:=(\mathbf{\Phi}(t))^{-1}\mathbf{X}_{t}=e^{-\mathbf{A}t}\mathbf{X}_{t}$ (note $\mathbf{Y}_{0}=\mathbf{X}_{0}$).
$d\mathbf{Y}_{t}=d\big(e^{-\mathbf{A}t}\mathbf{X}_{t}\big)=e^{-\mathbf{A}t}\,d\mathbf{X}_{t}+d(e^{-\mathbf{A}t})\,\mathbf{X}_{t}=$
$e^{-\mathbf{A}t}\big[(\mathbf{A}\mathbf{X}_{t}+\mathbf{b})\,dt+\sqrt{\epsilon}\,d\mathbf{W}_{t}\big]-\mathbf{A}e^{-\mathbf{A}t}\mathbf{X}_{t}\,dt=e^{-\mathbf{A}t}\mathbf{b}\,dt+\sqrt{\epsilon}\,e^{-\mathbf{A}t}d\mathbf{W}_{t},$
In the integral form:
$\mathbf{Y}_{t}=\mathbf{X}_{0}+\int_{0}^{t}e^{-\mathbf{A}s}\mathbf{b}\,ds+\sqrt{\epsilon}\int_{0}^{t}e^{-\mathbf{A}s}\,d\mathbf{W}_{s}.$
Multiplying by $\mathbf{\Phi}(t)=e^{\mathbf{A}t}$ yields:
$\mathbf{X}_{t}=e^{\mathbf{A}t}\mathbf{X}_{0}+\int_{0}^{t}e^{\mathbf{A}(t-s)}\mathbf{b}ds+\sqrt{\epsilon}\int_{0}^{t}e^{\mathbf{A}(t-s)}d\mathbf{W}_{s}.$ (15)
Hence, the conditional mean is:
$\bm{\mu}_{t|0}(\mathbf{X}_{0})=e^{\mathbf{A}t}\mathbf{X}_{0}+\Big(\!\int_{0}^{t}e^{\mathbf{A}(t-s)}ds\Big)\mathbf{b}=e^{\mathbf{A}t}X_{0}+\big(e^{\mathbf{A}t}-I\big)\mathbf{A}^{-1}\mathbf{b},$ (16)
Conditional variance.
$\mathbf{I}_{t}:=\int_{0}^{t}e^{A(t-s)}\,d\mathbf{W}_{s},\qquad\text{so that}\qquad\mathbf{X}_{t}-\bm{\mu}_{t|0}(\mathbf{X}_{0})=\sqrt{\epsilon}\,\mathbf{I}_{t}.$
$\mathrm{Cov}(\mathbf{X}_{t}\mid\mathbf{X}_{0})=\mathbb{E}\big[(\mathbf{X}_{t}-\bm{\mu}_{t|0})(\mathbf{X}_{t}-\bm{\mu}_{t|0})^{\top}\mid\mathbf{X}_{0}\big]=\epsilon\,\mathrm{Cov}(\mathbf{I}_{t}).$
In turn:
$\mathrm{Cov}(\mathbf{I}_{t})=\mathbb{E}\!\left[\left(\int_{0}^{t}e^{\mathbf{A}(t-s)}\,d\mathbf{W}_{s}\right)\left(\int_{0}^{t}e^{\mathbf{A}(t-r)}\,d\mathbf{W}_{r}\right)^{\!\top}\right].$
By Itô isometry:
$\mathbb{E}\!\left[\int_{0}^{t}G_{s}\,d\mathbf{W}_{s}\right]=0,\qquad\mathrm{Cov}\!\left(\int_{0}^{t}G_{s}\,d\mathbf{W}_{s},\int_{0}^{t}H_{s}\,dW_{s}\right)=\int_{0}^{t}G_{s}H_{s}^{\top}\,ds.$
Taking $G_{s}=H_{s}=e^{\mathbf{A}(t-s)}$ gives
$\mathrm{Cov}(\mathbf{I}_{t})=\int_{0}^{t}e^{\mathbf{A}(t-s)}\,e^{\mathbf{A}^{\top}(t-s)}\,ds.$
Because $\mathbf{A}$ is symmetric, $e^{\mathbf{A}^{\top}u}=e^{\mathbf{A}u}$, hence
$\mathrm{Cov}(\mathbf{I}_{t})=\int_{0}^{t}e^{2\mathbf{A}(t-s)}\,ds=\frac{1}{2}\big(e^{2\mathbf{A}t}-I\big)\mathbf{A}^{-1}.$
Consequently,
$\bm{\Sigma}_{t|0}=\mathrm{Cov}(\mathbf{X}_{t}\mid\mathbf{X}_{0})=\frac{\epsilon}{2}\big(e^{2\mathbf{A}t}-\mathbf{I}\big)\mathbf{A}^{-1}.$
Since $\mathbf{X}_{t}\mid\mathbf{X}_{0}$ is Gaussian with mean $\mu_{t|0}$ and covariance $\Sigma_{t|0}$, its score is
$\nabla_{\mathbf{X}_{t}}\log q(\mathbf{X}_{t}\mid\mathbf{X}_{0})=-\,\bm{\Sigma}_{t|0}^{-1}\big(\mathbf{X}_{t}-\bm{\mu}_{t|0}(\mathbf{X}_{0})\big).$
This completes the proof. ∎
Proof of Proposition 2.
Step 1: Joint law from the prior. From equation 15 and equation 16 in the proof of Proposition 1:
$\mathbf{X}_{u}=\bm{\mu}_{u|0}(\mathbf{X}_{0})+\sqrt{\epsilon}\int_{0}^{u}e^{\mathbf{A}(u-s)}d\mathbf{W}_{s},$
$\bm{\mu}_{u|0}(\mathbf{X}_{0})=e^{\mathbf{A}u}\mathbf{X}_{0}+\big(e^{\mathbf{A}u}-\mathbf{I}\big)\mathbf{A}^{-1}\mathbf{b}.$
Thus, conditionally on $\mathbf{X}_{0}$,
$\mathbb{E}[\mathbf{X}_{t}\mid\mathbf{X}_{0}]=\bm{\mu}_{t|0}(\mathbf{X}_{0}),\qquad\mathbb{E}[\mathbf{X}_{t^{\prime}}\mid\mathbf{X}_{0}]=\bm{\mu}_{t^{\prime}|0}(\mathbf{X}_{0}),$
$\bm{\Sigma}_{t|0}=\frac{\epsilon}{2}\big(e^{2\mathbf{A}t}-\mathbf{I}\big)\mathbf{A}^{-1},\qquad\bm{\Sigma}_{t^{\prime}|0}=\frac{\epsilon}{2}\big(e^{2\mathbf{A}t^{\prime}}-\mathbf{I}\big)\mathbf{A}^{-1}.$
For the cross-covariance, using Itô isometry and independence of increments, for $t<t^{\prime}$,
$\bm{\Sigma}_{t,t^{\prime}|0}=\mathrm{Cov}(\mathbf{X}_{t},\mathbf{X}_{t^{\prime}}\mid\mathbf{X}_{0})=\epsilon\int_{0}^{t}e^{\mathbf{A}(t-s)}\,e^{\mathbf{A}(t^{\prime}-s)}\,ds=\epsilon\int_{0}^{t}e^{\mathbf{A}(t+t^{\prime}-2s)}\,ds=\frac{\epsilon}{2}\,\mathbf{A}^{-1}\!\left(e^{\mathbf{A}(t+t^{\prime})}-e^{\mathbf{A}(t^{\prime}-t)}\right).$
Collecting blocks, we have the joint Gaussian (conditionally on $X_{0}$)
[…]
Step 2: Conditioning to obtain the bridge. For a joint Gaussian $\begin{bmatrix}x\\ y\end{bmatrix}$ with blocks $(\mu_{x},\mu_{y},\Sigma_{xx},\Sigma_{yy},\Sigma_{xy})$, the conditional $x\mid y$ is (Petersen et al., 2008, Section 8.1.3):
$x\mid y\sim\mathcal{N}\big(\mu_{x}+\Sigma_{xy}\Sigma_{yy}^{-1}(y-\mu_{y}),\Sigma_{xx}-\Sigma_{xy}\Sigma_{yy}^{-1}\Sigma_{yx}\big).$
Applying this with $x=\mathbf{X}_{t}$, $y=\mathbf{X}_{t^{\prime}}$ and the blocks above gives:
$\bm{\mu}_{t\mid 0,t^{\prime}}=\bm{\mu}_{t|0}(\mathbf{X}_{0})+\bm{\Sigma}_{t,t^{\prime}|0}\,\bm{\Sigma}_{t^{\prime}|0}^{-1}\!\big(\mathbf{X}_{t^{\prime}}-\bm{\mu}_{t^{\prime}|0}(\mathbf{X}_{0})\big),$
$\bm{\Sigma}_{t\mid 0,t^{\prime}}=\bm{\Sigma}_{t|0}-\bm{\Sigma}_{t,t^{\prime}|0}\,\bm{\Sigma}_{t^{\prime}|0}^{-1}\bm{\Sigma}_{t,t^{\prime}|0}.$
Here
$\bm{\Sigma}_{t|0}=\tfrac{\epsilon}{2}(e^{2\mathbf{A}t}-\mathbf{I})\mathbf{A}^{-1},\quad\bm{\Sigma}_{t^{\prime}|0}=\tfrac{\epsilon}{2}(e^{2\mathbf{A}t^{\prime}}-\mathbf{I})\mathbf{A}^{-1},\quad\bm{\Sigma}_{t,t^{\prime}|0}=\tfrac{\epsilon}{2}A^{-1}\big(e^{\mathbf{A}(t+t^{\prime})}-e^{\mathbf{A}(t^{\prime}-t)}\big).$
This completes the proof. ∎
Proof of Proposition 3.
Consider the following bijective reparameterization:
$v_{\phi}(\mathbf{X}_{t},t)=-\bm{\Sigma}_{t|0}^{-1}\left(\mathbf{X}_{t}-\bm{\mu}_{t|0}(\widehat{\mathbf{X}}_{0}^{\phi}(\mathbf{X}_{t},t))\right)$
and substitute it in the optimization problem:
$\min_{\phi}\mathbb{E}_{\mathbf{X}_{0},\mathbf{X}_{t},t}\left[\left\|v_{\phi}(\mathbf{X}_{t},t)+\bm{\Sigma}_{t|0}^{-1}(\mathbf{X}_{t}-\bm{\mu}_{t|0}(\mathbf{X}_{0}))\right\|^{2}\right],$
$\min_{\phi}\mathbb{E}_{\mathbf{X}_{0},\mathbf{X}_{t},t}\left[\left\|-\bm{\Sigma}_{t|0}^{-1}\left(\mathbf{X}_{t}-\bm{\mu}_{t|0}(\widehat{\mathbf{X}}_{0}^{\phi}(\mathbf{X}_{t},t))\right)+\bm{\Sigma}_{t|0}^{-1}(\mathbf{X}_{t}-\bm{\mu}_{t|0}(\mathbf{X}_{0}))\right\|^{2}\right]=$
$\min_{\phi}\mathbb{E}_{\mathbf{X}_{0},\mathbf{X}_{t},t}\left[\left\|\bm{\Sigma}_{t|0}^{-1}\left((\mathbf{X}_{t}-\bm{\mu}_{t|0}(\mathbf{X}_{0}))-\left(\mathbf{X}_{t}-\bm{\mu}_{t|0}(\widehat{\mathbf{X}}_{0}^{\phi}(\mathbf{X}_{t},t))\right)\right)\right\|^{2}\right]=$
$\min_{\phi}\mathbb{E}_{\mathbf{X}_{0},\mathbf{X}_{t},t}\left[\left\|\bm{\Sigma}_{t|0}^{-1}\left(\bm{\mu}_{t|0}(\widehat{\mathbf{X}}_{0}^{\phi}(\mathbf{X}_{t},t))-\bm{\mu}_{t|0}(\mathbf{X}_{0})\right)\right\|^{2}\right]=$
$\min_{\phi}\mathbb{E}_{\mathbf{X}_{0},\mathbf{X}_{t},t}\left[\left(\bm{\mu}_{t|0}(\widehat{\mathbf{X}}_{0}^{\phi}(\mathbf{X}_{t},t))-\bm{\mu}_{t|0}(\mathbf{X}_{0})\right)^{\top}(\bm{\Sigma}_{t|0}^{\top})^{-1}\bm{\Sigma}_{t|0}^{-1}\left(\bm{\mu}_{t|0}(\widehat{\mathbf{X}}_{0}^{\phi}(\mathbf{X}_{t},t))-\bm{\mu}_{t|0}(\mathbf{X}_{0})\right)\right]$
Taking the gradient of this objective with respect to $\phi$, we obtain:
$\mathbb{E}_{\mathbf{X}_{0},\mathbf{X}_{t},t}\left[2(\bm{\Sigma}_{t|0}^{\top})^{-1}\bm{\Sigma}_{t|0}^{-1}\left(\bm{\mu}_{t|0}(\widehat{\mathbf{X}}_{0}^{\phi}(\mathbf{X}_{t},t))-\bm{\mu}_{t|0}(\mathbf{X}_{0})\right)\right]=0$
Since $\bm{\Sigma}_{t|0}^{-1}$ is positive definite we can multiply by $\frac{1}{2}\bm{\Sigma}_{t|0}(\bm{\Sigma}_{t|0}^{\top})$ and get:
$\mathbb{E}_{\mathbf{X}_{0},\mathbf{X}_{t},t}\left[\bm{\mu}_{t|0}(\widehat{\mathbf{X}}_{0}^{\phi}(\mathbf{X}_{t},t))-\bm{\mu}_{t|0}(\mathbf{X}_{0})\right]=0$
Then for each $\mathbf{X}_{t}$ consider conditional mean:
$\mathbb{E}_{\mathbf{X}_{t},t}\left[\mathbb{E}_{\mathbf{X}_{0}|\mathbf{X}_{t}}\left[\bm{\mu}_{t|0}(\widehat{\mathbf{X}}_{0}^{\phi}(\mathbf{X}_{t},t))-\bm{\mu}_{t|0}(\mathbf{X}_{0})\right]\right]=0$
$\mathbb{E}_{\mathbf{X}_{0}|\mathbf{X}_{t}}\left[\bm{\mu}_{t|0}(\widehat{\mathbf{X}}_{0}^{\phi}(\mathbf{X}_{t},t))-\bm{\mu}_{t|0}(\mathbf{X}_{0})\right]=0$
From equation 16 we have:
$\mu_{t|0}(\mathbf{X}_{0})=e^{\mathbf{A}t}\mathbf{X}_{0}+\Big(\!\int_{0}^{t}e^{\mathbf{A}(t-s)}ds\Big)\mathbf{b}=e^{\mathbf{A}t}X_{0}+\big(e^{\mathbf{A}t}-I\big)\mathbf{A}^{-1}\mathbf{b},$
Then (note that $e^{\mathbf{A}t}$ is invertible and we can multiplu both sides on $e^{-At}$):
$\mathbb{E}_{\mathbf{X}_{0}|\mathbf{X}_{t}}\left[e^{\mathbf{A}t}\mathbf{X}_{0}^{\phi}(\mathbf{X}_{t},t)+\big(e^{\mathbf{A}t}-I\big)\mathbf{A}^{-1}\mathbf{b}-e^{\mathbf{A}t}\mathbf{X}_{0}+\big(e^{\mathbf{A}t}-I\big)\mathbf{A}^{-1}\mathbf{b}\right]=0$
$\mathbb{E}_{\mathbf{X}_{0}|\mathbf{X}_{t}}\left[e^{\mathbf{A}t}(\mathbf{X}_{0}^{\phi}(\mathbf{X}_{t},t)-\mathbf{X}_{0})\right]=0$
$\mathbb{E}_{\mathbf{X}_{0}|\mathbf{X}_{t}}\left[\mathbf{X}_{0}^{\phi}(\mathbf{X}_{t},t)-\mathbf{X}_{0}\right]=0$
$\mathbf{X}_{0}^{\phi}(\mathbf{X}_{t},t)=\mathbb{E}_{\mathbf{X}_{0}|\mathbf{X}_{t}}\left[\mathbf{X}_{0}\right]$
Hence, optimal $\mathbf{X}_{0}^{*}=\mathbb{E}_{\mathbf{X}_{0}|\mathbf{X}_{t}}\left[\mathbf{X}_{0}\right]$, which in turn is the minimizer of MSE problem:
$\min_{\phi}\mathbb{E}_{\mathbf{X}_{0},\mathbf{X}_{t},t}\left[\|\widehat{\mathbf{X}}_{0}^{\phi}(\mathbf{X}_{t},t)-\mathbf{X}_{0}\|^{2}\right].$
By substituting in to $v_{\phi}$ we have:
$v^{*}(\mathbf{X}_{t},t)=-\bm{\Sigma}_{t|0}^{-1}\left(\mathbf{X}_{t}-\bm{\mu}_{t|0}(\widehat{\mathbf{X}}_{0}^{*}(\mathbf{X}_{t},t))\right)$
This completes the proof. ∎
Appendix B Implementation Details
To evaluate the performance of our method and compare it to other approaches, we take our own small and simple U-Net model based on several residual blocks with $2D$ convolutions. The model has approximately $8.7$ million parameters. For each experiment, we extract sub-sequences from the videos, consisting of $10$ consecutive frames, and concatenate them into a $10$-channel input for the neural network. We use $9,500$ training sequences and $500$ validation sequences, with a random split for each seed value. We train each model for $150,000$ iterations with a batch size of $128$ and an ema rate of $0.999$. We use the AdamW optimizer (Loshchilov and Hutter, 2019) with betas set to $0.9$ and $0.95$, a weight decay of $10^{-4}$, and a learning rate of $3\times 10^{-5}$. The number of steps in the reverse process for all methods is equal to $1000$. All experiments are conducted using a single NVIDIA Tesla A100100 GPU.
Appendix C Additional Quantitative Results
Here, we present the results of a quantitative comparison between DDPM, DDIM, Bridge Matching (BM), and our proposed method, TCVBM. The values of the standard deviation are provided, based on 3 runs of each method with different random seeds.
Table 4: Frame interpolation quantitative results with standard deviation. The best values in column are bold, second best values are underlined.
Metric FVD $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$
DDIM 34.664 $\pm$ 5.80 0.120 $\pm$ 0.070 15.843 $\pm$ 0.120 0.766 $\pm$ 0.011
DDPM 33.612 $\pm$ 1.494 0.107 $\pm$ 0.009 14.509 $\pm$ 0.427 0.714 $\pm$ 0.024
BM 34.766 $\pm$ 0.398 0.078 $\pm$ 0.001 17.265 $\pm$ 0.390 0.789 $\pm$ 0.005
TCVBM (ours) 31.491 $\pm$ 4.035 0.071 $\pm$ 0.019 17.451 $\pm$ 0.459 0.825 $\pm$ 0.044
Table 5: Image-to-Video generation quantitative results with standard deviation. The best values in column are bold, second best values are underlined.
Metric FVD $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$
DDIM 335.51 $\pm$ 241.12 0.402 $\pm$ 0.092 10.205 $\pm$ 0.514 0.513 $\pm$ 0.069
DDPM 250.52 $\pm$ 134.99 0.383 $\pm$ 0.054 10.333 $\pm$ 0.275 0.530 $\pm$ 0.046
BM 48.54 $\pm$ 0.56 0.268 $\pm$ 0.005 10.627 $\pm$ 0.053 0.582 $\pm$ 0.004
TCVBM (ours) 45.32 $\pm$ 0.91 0.260 $\pm$ 0.002 10.710 $\pm$ 0.028 0.589 $\pm$ 0.001
Table 6: Video super resolution quantitative results with standard deviation. The best values in column are bold, second best values are underlined.
Metric FVD $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$
DDIM 336.808 $\pm$ 5.175 0.520 $\pm$ 0.006 17.226 $\pm$ 0.103 0.600 $\pm$ 0.016
DDPM 614.288 $\pm$ 6.289 0.237 $\pm$ 0.001 20.152 $\pm$ 0.050 0.577 $\pm$ 0.004
BM 29.710 $\pm$ 20.683 0.026 $\pm$ 0.005 21.412 $\pm$ 1.040 0.941 $\pm$ 0.012
TCVBM (ours) 32.762 $\pm$ 23.153 0.029 $\pm$ 0.004 21.431 $\pm$ 1.419 0.941 $\pm$ 0.011
Appendix D Initialization Experiments
In this section, we explore options for initializing or representing input data for bridge-based methods used in our work, namely for Bridge Matching with Browninan Bridge (BM) and Time-Correlated Video Bridge Matching (TCVBM). The interest in exploring the effect of input data initialization on the quality of model performance stems primarily from the assumption that bridge-based approaches are better suited for data-to-data translation tasks.
D.1 Frame Interpolation
As input data for the network, we explored two options: filling in intermediate frames with Gaussian noise sampled from $\mathcal{N}(\textbf{0},\textbf{1})$ and using linear interpolation between fixed boundary frames $\mathbf{x}^{0}$ and $x^{N}$, i.e.:
$\mathbf{x}^{n}_{input}=\frac{n\mathbf{x}^{0}+(N-n)x^{N}}{N},\quad n=1,\dots,N-1.$
Table 7 compares the results of these initialization methods. As can be seen, filling intermediate frames with noise from a normal distribution produces better results than the initial linear interpolation.
Table 7: Analysis of the impact of initialization of input video data for bridge-based methods in the task of frame interpolation. Using noise from a normal distribution shows a clear advantage. The best values in column are bold, second best values are underlined.
Initialization method Method FVD $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$
Linear interpolation BM 34.804 0.109 15.439 0.756
TCVBM 31.944 0.092 16.275 0.782
Gaussian noise from $\mathcal{N}(\textbf{0},\textbf{1})$ BM 34.315 0.079 17.103 0.794
TCVBM 30.542 0.077 17.280 0.813
D.2 Image-to-Video Generation
Here we compare the following two types of initial initialization: duplicating the first frame in place of the frames to be generated (static video) and using random noise everywhere except the first frame. Static video initialization is superior to the noise option for both models (Table 8).
Table 8: Comparison of two types of initial initialization of input data for image-to-video generation.
Initialization method Method FVD $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$
Static video BM 49.32 0.271 10.63 0.579
TCVBM 44.96 0.258 10.75 0.591
Gaussian noise from $\mathcal{N}(\textbf{0},\textbf{1})$ BM 52.57 0.287 10.61 0.568
TCVBM 48.61 0.263 10.68 0.587
D.3 Video Super Resolution
Table 9: Comparison of two types of initial initialization of input data for video super resolution. We perform this comparison for low-resolution $32\times 32$.
Initialization method Method FVD $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$
Low-resolution video BM 9.501 0.014 24.888 0.972
TCVBM 9.496 0.012 24.970 0.973
Low-resolution video BM 9.556 0.012 24.892 0.973
concatenated with noise from $\mathcal{N}(\textbf{0},\textbf{1})$ TCVBM 9.646 0.012 24.988 0.973
Appendix E Hyperparameters Searching
Here we compare different values of hyperparameters, namely the noise scaling value $\epsilon$ and the coefficient $\alpha$, which determines the degree of impact of the matrix $\mathbf{A}$ as $\widetilde{\mathbf{A}}:=\alpha\mathbf{A}$. Tables 10 and  11 show that in the case of frame interpolation and image-to-video generation, it is impossible to identify a clear dependence of the generation quality on the hyperparameters used, however, a sufficient amount of noise and not large values for the $\alpha$ coefficient are optimal. The results for video super resolution in Table 12 demonstrate that small values of $\epsilon$ and $\alpha$ are optimal for this task, which does not contradict the results for frame interpolation.
Table 10: The results of TCVBM training with various hyperparameters $\epsilon$ and $\alpha$ for frame interpolation. The best values in column are bold, second best values are underlined.
$\epsilon$ $\alpha$ FVD $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$
0.1 0.1 35.572 0.085 16.40 0.797
0.1 1 36.542 0.089 16.37 0.797
0.1 10 29.792 0.086 16.56 0.801
1 0.1 27.342 0.084 16.65 0.803
1 1 30.542 0.077 16.86 0.813
1 10 31.432 0.080 17.12 0.817
10 0.1 37.662 0.084 17.24 0.819
10 1 54.542 0.093 17.17 0.818
10 10 54.562 0.100 17.17 0.769
Table 11: The results of TCVBM training with various hyperparameters $\epsilon$ and $\alpha$ for image-to-video generation. The best values in column are bold, second best values are underlined.
$\epsilon$ $\alpha$ FVD $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$
0.1 0.1 55.46 0.2670 10.80 0.587
0.1 1 51.62 0.2578 10.74 0.583
0.1 10 59.57 0.2571 10.67 0.585
1 0.1 51.80 0.2615 10.70 0.590
1 1 44.96 0.2575 10.75 0.591
1 10 51.27 0.2613 10.60 0.583
10 0.1 44.90 0.2610 10.67 0.588
10 1 44.22 0.2584 10.70 0.591
10 10 44.58 0.2597 10.58 0.583
Table 12: The results of TCVBM training with various hyperparameters $\epsilon$ and $\alpha$ for video super resolution. We perform this comparison for low-resolution $32\times 32$.
$\epsilon$ $\alpha$ FVD $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$
0.1 0.1 9.496 0.012 24.970 0.973
0.1 1 10.413 0.013 24.358 0.969
1 0.1 13.226 0.019 23.004 0.959
1 1 15.023 0.022 22.120 0.949
10 0.1 18.458 0.033 20.085 0.921
10 1 20.746 0.039 19.283 0.906
Appendix F Computational Cost Analysis
As described in Algorithm 2, inference consists of two alternating steps: (i) computing $\widehat{\mathbf{X}}_{0}^{\phi}(\mathbf{X}_{t_{n}},t_{n})$ and (ii) sampling from the Gaussian distribution given in Eq. 9. Sampling can be done using reparameterization:
$\mathbf{X}_{t^{\prime}}=\mathbf{\mu}_{t|0,t^{\prime}}(\mathbf{X}_{0})+(\mathbf{\Sigma}_{t|0,t^{\prime}}^{1/2})^{\top}\mathbf{Z},\quad\mathbf{Z}\sim\mathcal{N}(0,I),$
which requires the evaluation of the mean and covariance:
$\mathbf{\mu}_{t|0,t^{\prime}}=\mathbf{\mu}_{t|0}(\mathbf{X}_{0})+\mathbf{\Sigma}_{t|0}\mathbf{\Sigma}_{t^{\prime}|0}^{-1}(\mathbf{X}_{t^{\prime}}-\mathbf{\mu}_{t^{\prime}|0}(\mathbf{X}_{0})),\quad$
$\mathbf{\Sigma}_{t|0,t^{\prime}}=\mathbf{\Sigma}_{t|0}-\mathbf{\Sigma}_{t|0}\mathbf{\Sigma}_{t^{\prime}|0}^{-1}\mathbf{\Sigma}_{t|0}.$
In turn,
$\mathbf{\mu}_{t|0}(\mathbf{X}_{0})=e^{\mathbf{A}t}\mathbf{X}_{0}+\left(e^{\mathbf{A}t}-I\right)\mathbf{A}^{-1}\mathbf{b},\quad\mathbf{\Sigma}_{t|0}=\epsilon\frac{e^{2\mathbf{A}t}-I}{2}\mathbf{A}^{-1}.$
In total, almost all these operations can be cached except for 5 matrix multiplications: $(\mathbf{\Sigma}_{t|0,t^{\prime}}^{1/2})^{\top}\mathbf{Z}$, $e^{\mathbf{A}t}\mathbf{X_{0}}$, $(\mathbf{\Sigma}_{t|0}\mathbf{\Sigma}_{t^{\prime}|0}^{-1})\mathbf{X}_{t^{\prime}}$, $e^{\mathbf{A}t^{\prime}}\mathbf{X_{0}}$, $(\mathbf{\Sigma}_{t|0}\mathbf{\Sigma}_{t^{\prime}|0}^{-1})(e^{\mathbf{A}t^{\prime}}\mathbf{X_{0}})$. All these multiplications are for tensors of size $F\times F$ ($F$ is the number of frames in the video) with a tensor of size $F\times C\times H\times W$, where $C$ is the number of channels, $H$ is the height, and $W$ is the width of the video. This requires $O(F^{2}CHW)$ operations. For comparison, a single convolutional layer with $C$ input channels, $C_{\text{out}}$ output channels, and kernel size $K\times K$ applied to all $F$ frames requires $O(FCHWK^{2}C_{\text{out}})$ operations. Since typically $F\sim 10{-}100$, $K^{2}\sim 10$, and $C_{\text{out}}\sim 10$, we have
$\frac{F^{2}CHW}{FCHWK^{2}C_{\text{out}}}=\frac{F}{K^{2}C_{\text{out}}}\sim 0.1-1.$
Thus, each of these 5 operations is comparable to requiring less computation than 1 convolution layer. Since a neural network requires many convolutional layers and additional nonlinear processing, the resulting overhead of our bridge update is comparable to only a few convolutional layers and therefore remains relatively small in practice.
Appendix G Dynamical Correlation
G.1 Theory
Consider prior SDE with an additional multiplicative function $f(t)$ depending only on time $t$:
$d\mathbf{X}_{t}=f(t)(\mathbf{A}\mathbf{X}_{t}+\mathbf{b})dt+\sqrt{\epsilon}d\mathbf{W}_{t}.$
The derivation of formulas for this prior follows the same principles used for $f(t)=1$ in Appendix A.
Define:
$F(t)=\int_{0}^{t}f(\tau)\,d\tau,\qquad\mathbf{\Phi}(t)=e^{\mathbf{A}\,F(t)}.$
Then $\frac{d}{dt}\mathbf{\Phi}(t)=f(t)\,\mathbf{A}\,\mathbf{\Phi}(t)$, $\mathbf{\Phi}(0)=I$, and $\frac{d}{dt}\mathbf{\Phi}(t)^{-1}=-\,f(t)\,\mathbf{\Phi}(t)^{-1}\mathbf{A}$.
Conditional mean. Consider $\mathbf{Y}_{t}:=\mathbf{\Phi}(t)^{-1}\mathbf{X}_{t}$. By Itô’s rule:
$d\mathbf{Y}_{t}=\mathbf{\Phi}(t)^{-1}d\mathbf{X}_{t}+d(\mathbf{\Phi}(t)^{-1})\,\mathbf{X}_{t}=\mathbf{\Phi}(t)^{-1}f(t)(\mathbf{A}\mathbf{X}_{t}+\mathbf{b})\,dt+\sqrt{\epsilon}\,\mathbf{\Phi}(t)^{-1}d\mathbf{W}_{t}-f(t)\,\mathbf{\Phi}(t)^{-1}A\mathbf{X}_{t}\,dt=$
$=\mathbf{\Phi}(t)^{-1}f(t)\mathbf{b}\,dt+\sqrt{\epsilon}\,\mathbf{\Phi}(t)^{-1}d\mathbf{W}_{t}.$
Integrating from 0 to $t$ gives
$\mathbf{Y}_{t}=\mathbf{X}_{0}+\int_{0}^{t}\mathbf{\Phi}(s)^{-1}f(s)\,\mathbf{b}\,ds+\sqrt{\epsilon}\int_{0}^{t}\mathbf{\Phi}(s)^{-1}d\mathbf{W}_{s},$
and thus
$\mathbf{X}_{t}=\mathbf{\Phi}(t)\mathbf{X}_{0}+\mathbf{\Phi}(t)\!\int_{0}^{t}\mathbf{\Phi}(s)^{-1}f(s)\,\mathbf{b}\,ds+\sqrt{\epsilon}\,\mathbf{\Phi}(t)\!\int_{0}^{t}\mathbf{\Phi}(s)^{-1}d\mathbf{W}_{s}.$
Taking expectation and using $\mathbf{\Phi}(t)\mathbf{\Phi}(s)^{-1}=e^{\mathbf{A}(F(t)-F(s))}$,
$\bm{\mu}_{t|0}(\mathbf{X}_{0})=\mathbb{E}[\mathbf{X}_{t}|\mathbf{X}_{0}]=e^{\mathbf{A}F(t)}\mathbf{X}_{0}+\int_{0}^{t}e^{\mathbf{A}(F(t)-F(s))}f(s)\,\mathbf{b}\,ds.$
With the change of variables $u=F(s)$ (so du=f(s)dsdu=f(s)\,ds) this equals
$\int_{0}^{F(t)}e^{\mathbf{A}(F(t)-u)}\,du\,\mathbf{b}=\Big[-\,e^{\mathbf{A}(F(t)-u)}\mathbf{A}^{-1}\Big]_{u=0}^{F(t)}\mathbf{b}=\big(e^{\mathbf{A}F(t)}-I\big)\mathbf{A}^{-1}\mathbf{b}.$
Therefore
$\bm{\mu}_{t|0}(\mathbf{X}_{0})=e^{\mathbf{A}F(t)}\mathbf{X}_{0}+\big(e^{\mathbf{A}F(t)}-I\big)\mathbf{A}^{-1}\mathbf{b}.$
The mean $\mu_{t|0}(X_{0})$ of the process starting from a given $\mathbf{X}_{0}$ is given by:
$\bm{\mu}_{t|0}(\mathbf{X}_{0})=e^{\mathbf{A}F(t)}\mathbf{X}_{0}+\left(e^{\mathbf{A}F(t)}-I\right)\mathbf{A}^{-1}\mathbf{b}.$
Conditional variance. Let $\mathbf{Z}_{t}:=\mathbf{X}_{t}-\bm{\mu}_{t|0}$ be the centered process. Subtracting the mean SDE from the original SDE yields
$d\mathbf{Z}_{t}=f(t)\,\mathbf{A}\,\mathbf{Z}_{t}\,dt+\sqrt{\epsilon}\,d\mathbf{W}_{t}.$
Define the covariance $\bm{\Sigma}_{t|0}:=\mathbb{E}[\mathbf{Z}_{t}\mathbf{Z}_{t}^{\top}]$. Using Itô’s rule for $\mathbf{Z}_{t}\mathbf{Z}_{t}^{\top}$,
$d(\mathbf{Z}_{t}\mathbf{Z}_{t}^{\top})=(d\mathbf{Z}_{t})\mathbf{Z}_{t}^{\top}+\mathbf{Z}_{t}(d\mathbf{Z}_{t})^{\top}+d\mathbf{Z}_{t}(d\mathbf{Z}_{t})^{\top},$
where
$(d\mathbf{Z}_{t})\,\mathbf{Z}_{t}^{\top}=f(t)\mathbf{A}\mathbf{Z}_{t}\mathbf{Z}_{t}^{\top}\,dt+\sqrt{\epsilon}\,d\mathbf{W}_{t}\,\mathbf{Z}_{t}^{\top},$ (17)
$\mathbf{Z}_{t}\,(d\mathbf{Z}_{t})^{\top}=f(t)\mathbf{Z}_{t}\mathbf{Z}_{t}^{\top}\mathbf{A}^{\top}\,dt+\sqrt{\epsilon}\,\mathbf{Z}_{t}\,d\mathbf{W}_{t}^{\top},$ (18)
$(d\mathbf{Z}_{t})(d\mathbf{Z}_{t})^{\top}=\epsilon\,d\mathbf{W}_{t}d\mathbf{W}_{t}^{\top}=\epsilon\,I\,dt.$ (19)
Hence
$d(\mathbf{Z}_{t}\mathbf{Z}_{t}^{\top})=f(t)\big(\mathbf{A}\mathbf{Z}_{t}\mathbf{Z}_{t}^{\top}+\mathbf{Z}_{t}\mathbf{Z}_{t}^{\top}\mathbf{A}^{\top}\big)\,dt+\sqrt{\epsilon}\,d\mathbf{W}_{t}\,\mathbf{Z}_{t}^{\top}+\sqrt{\epsilon}\,\mathbf{Z}_{t}\,d\mathbf{W}_{t}^{\top}+\epsilon\,I\,dt.$
and taking expectations gives
$\frac{d}{dt}\bm{\Sigma}_{t|0}=f(t)\,\mathbf{A}\,\bm{\Sigma}_{t|0}+f(t)\,\bm{\Sigma}_{t|0}\mathbf{A}^{\top}+\epsilon\,I,\quad\bm{\Sigma}_{t|0}=0.$
To get the cross-covariance, we use:
$\mathbf{Z}_{t}=\sqrt{\epsilon}\int_{0}^{t}e^{\mathbf{A}(F(t)-F(s))}\,d\mathbf{W}_{s},\qquad\mathbf{Z}_{t^{\prime}}=\sqrt{\epsilon}\int_{0}^{t^{\prime}}e^{\mathbf{A}(F(t^{\prime})-F(u))}\,d\mathbf{W}_{u},$
The Itô isometry yields
$\bm{\Sigma}_{t,t^{\prime}|0}:=\mathrm{Cov}(\mathbf{Z}_{t},\mathbf{Z}_{t^{\prime}}^{\top})=\epsilon\int_{0}^{t}e^{\mathbf{A}(F(t)-F(s))}\,e^{\mathbf{A}(F(t^{\prime})-F(s))^{\top}}\,ds.$
For symmetric $\mathbf{A}$, $e^{(\cdot)^{\top}}=e^{(\cdot)}$ and, since these exponentials commute (all are $e^{\mathbf{A}(\cdot)}$),
$e^{\mathbf{A}(F(t)-F(s))}\,e^{\mathbf{A}(F(t^{\prime})-F(s))}=e^{\mathbf{A}(F(t^{\prime})-F(t))}\,e^{2\mathbf{A}(F(t)-F(s))}.$
Hence
$\bm{\Sigma}_{t,t^{\prime}|0}=e^{\mathbf{A}(F(t^{\prime})-F(t))}\,\epsilon\int_{0}^{t}e^{2\mathbf{A}(F(t)-F(s))}\,ds=e^{\mathbf{A}\,[F(t^{\prime})-F(t)]}\,\bm{\Sigma}_{t|0}.$
Summary. Thus, all three components: mean $\bm{\mu}_{t|0}(\mathbf{X}_{0})$, variance $\bm{\Sigma}_{t|0}$, and cross-covariance $\bm{\Sigma}_{t,t^{\prime}|0}$ are derived and can be used further in the same way as in the original case of $f(t)=1$.
G.2 Experimental Results
The continuous and time-decreasing function $f(t)$ sets the increasing values of the matrix $\mathbf{A}$ in the inverse diffusion process when $t\rightarrow 0$. Thus, the correlation of frames with each other in the generated video increases in the last steps of the inference. We conduct a series of experiments to investigate the effect of dynamic correlation and the choice of the function $f(t)$. We use the same experimental setting as described in Appendix B of the main paper, with a number of optimizer steps set to $120,000$. The frame interpolation results are presented in Table 13. As can be seen, our experiments do not demonstrate the advantages of using dynamic correlation compared to using constant values of the matrix $\mathbf{A}$. However, we observe significant differences in the quality of the results depending on the function $f(t)$. This demonstrates a complex structure of the relationship between video frames in the diffusion process, which our framework provides in the constant linear approximation.
Table 13: The results of the selection of the function $f(t)$, which determines the inverse dependence of the values of the matrix $\mathbf{A}$ on time $t$. The best values in column are bold, second best values are underlined.
$f(t)$ FVD $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$
1t1-t 49.569 0.115 13.938 0.752
12t1-2t 398.823 0.202 11.760 0.684
10.5t1-0.5t 427.942 0.226 12.063 0.620
$2\times(1-t)$ 409.495 0.195 12.239 0.624
$0.5\times(1-t)$ 269.899 0.182 12.297 0.676
$(1-t)^{2}$ 43.651 0.112 13.916 0.751
$e^{-t}$ 1910.002 0.973 3.600 0.002
$e^{-2t}$ 1216.250 0.629 7.243 0.119
$e^{-4t}$ 108.052 0.142 13.186 0.688
$e^{-8t}$ 50.078 0.128 13.321 0.726
Constant ($f(t)=\alpha=1$) 36.309 0.097 14.515 0.772
Appendix H Additional Examples
H.1 Frame Interpolation
Figure 3: Additional frame interpolation results.
H.2 Image-to-Video Generation
Figure 4: Additional image-to-video results.
H.3 Video Super Resolution
Figure 5: Additional video super resolution results. The resolution is increased from $16\times 16$ to $64\times 64$.